### v1

# Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior

Abstract
Animals flexibly recombine a finite set of core motor primitives to meet diverse task demands, but existing behavior-segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To reflect the animal behavior generation procedure, we introduce skill-based imitation learning (SKIL) for behavior understanding, a reinforcement learning-based imitation framework that (1) infers interpretable skill sets, i.e., latent basis functions of behavior, by leveraging representation learning on transition probabilities, and (2) parameterizes policies as dynamic mixtures of these skills. We validate our approach on a simple grid world, a discrete labyrinth, and unconstrained videos of freely moving animals. Across tasks, it identifies reusable skill components, learns continuously evolving compositional policies, and generates realistic trajectories beyond the capabilities of traditional discrete models. By exploiting generative behavior modeling with compositional representations, our method offers a concise, principled account of how complex animal behaviors emerge from dynamic combinations of fundamental motor primitives.
1 Introduction
Animal behavior is composed of recurring patterns that reflect underlying motor primitives like back grooming, running, and sniffing. Behavior segmentation methods [1, 2, 3, 4, 5] seek to uncover structured patterns in behavior by dividing continuous pose trajectories into discrete units known as stereotyped behavioral syllables. These segmented units can then be used to build structured representations of movement for downstream neurobehavioral analysis.
However, existing behavior segmentation methods face three major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often abstract actions into high-level syllables, overlooking shared motion components across body parts. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering [4, 5]) or rely on restrictive generative assumptions (e.g., linear dynamics and hidden Markov model [1, 2, 3]), often leading to unrealistic synthesized behaviors.
To address these limitations, we introduce a new perspective: modeling behavior as emerging from an agent following a reinforcement learning (RL) policy. We study behavioral dynamics and motor primitives by inferring the animal’s policy through an RL–based imitation learning framework. It not only enables more realistic behavior generation through RL but also allows us to discover motor primitives as reusable skill sets that drive behavior toward internal goals. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor primitives that go beyond the constraints of discrete segmentation.
Concretely, we first learn interpretable latent representations from behavior trajectories, which we refer to as skill sets. These represent low-level motion patterns that act as modular building blocks of behavior. Such skills can involve different body parts. By flexibly combining these skill sets, animals can generate a wide range of complex, high-level behaviors. For example, face grooming (forepaw-to-face movements) and body grooming (repetitive strokes along the torso) share common grooming-related skills, while also engaging distinct body parts specific to each movement. We leverage existing advances in representation learning in RL [6, 7, 8] to learn skill sets from transition probability functions. Then, we parametrize policies using these skill sets and apply imitation learning [9] to train the policies from animal behavior demonstrations.
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor skills.
Our method is referred to as SKIL, short for Skill-Based Imitation Learning. The fine-grained skill representation enables highly flexible behavioral characterization through the policy. Critically, over long behavioral trajectories, we assume an animal draws from a fixed set of core motor skills, regardless of the specific movements performed [10, 11]. The policy captures how the contribution of each skill evolves continuously over time, reflecting dynamic behavioral changes. Due to the spatial specificity of different motor components, some skills may be engaged briefly, while others persist over longer periods. Moreover, multiple skills can be active simultaneously to construct ongoing movement (Fig. 1). This flexible, compositional view of behavior cannot be achieved using traditional discrete behavioral syllables. SKIL thus offers a more nuanced understanding of how complex actions are generated from dynamic combinations of fundamental motor primitives. Such insight enables researchers to test whether specific fine-grained motor primitives depend on certain local neural circuits, offering a more precise understanding of movement control and its underlying neural mechanisms.
In the remainder of the paper, we introduce preliminaries in Sec. 2, which pave the way for our methods for animal behavior understanding in Sec. 3. We review related work in Sec. 4 and demonstrate the advantage of our methods empirically in Sec. 5. Finally, we conclude with a discussion in Sec. 6.
2 Preliminaries
Markov Decision Processes (MDP). To define skills that characterize mouse behavior, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and the animal’s condition—for example, as defined by pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that the animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
Following standard notations, we define the value function $V(s):=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the action-value function $Q(s,a)=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$, which are the expected discounted cumulative rewards when executing policy $\pi$. From the above definition, we can establish the following Bellman relationship:
$Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl{[}V_{h+1}(s^{\prime})\bigr{]},\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl{[}Q_{h}(s,a)\bigr{]}.$ (1)
Imitation Learning. The goal of imitation learning(IL) is to find a policy $\pi$ that mimics the behavior of the animal demonstrations. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
[…]
$f^{*}=\mathop{\arg\max}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\log\frac{f(x_{i},c_{i})}{f(x_{i},c_{i})+\sum_{j=1}^{k}f(y_{i,j},c_{i})},$ (2)
where $\{(x_{i},c_{i})\}_{i=1}^{n}$ are observed samples, $\{y_{i,j}\}_{i\in[n],j\in[k]}$ are artificially generated data sampled from a predefined noise distribution $q(y)$. Here, $k$ represents the ratio of negative to positive samples. This objective can be interpreted as a multi-class extension for distinguishing real data from noise.
3 Skill-based Imitation learning (SKIL) for Behavior Understanding
Given the MDP definition, we frame skill discovery from a control-theoretic perspective. In this view, skills are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 1 (Skill Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in a linear (or low-rank) MDP [15, 16, 17], we can express it via a spectral decomposition:
$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime}),$ (3)
where $\phi:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:{\mathcal{S}}\rightarrow\mathbb{R}^{d}$, and $q(s^{\prime})$ is some parametrized probability measure supported on the state space $\mathcal{S}$. We define the function $\phi$ as the skill set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Because animal behavior (e.g., navigation, pose dynamics) follows structured, stereotyped patterns that evolve on low-dimensional manifolds, it can be effectively approximated by low-rank MDPs even in the presence of underlying nonlinearities.
Given the skill definition, substituting Eq. (3) and the linearized reward model into the Bellman equation (Eq. (1)), we get:
$\!Q(s,a)$ $=r(s,a)+\gamma\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (4)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of skill features $\phi(s,a)$, offering a convenient way to link skills to the policy. Following the maximum entropy reinforcement learning framework [18], we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal policy $\pi(a|s)$ can be shown to follow:
$\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (5)
where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl{(}\pi(a|s)\bigr{)}$ is the entropy.
Thereby, the skill function $\phi(s,a)$ not only spans the space of $Q$-functions but also provides a universal basis for constructing the policy $\pi(a|s)$ via Eq. (5). Since $\phi(s,a)$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward function or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor skills available to the animal, while the weight vector $u$ captures task-specific modulations required to produce behavior consistent with different goals. With the skill set established as a foundational representation for both value functions and policies, we are equipped to interpret behavioral trajectories through the lens of skill alignment.
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (skill discovery), as well as $u$ (skill-based policy learning) from state-action-next-state observations. The learning procedure differs depending on the nature of the behavior data.
3.1 Discrete Version
Skill discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) [19, 20, 21] or spectral decomposition representation [7, 22] to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting skill set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Skill-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. (5), using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (6)
Limitations for continuous data. While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the skill discovery step (Eq. (3)), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals [2]. For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in skill-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. (6) is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the skill representations and policy in continuous state-action spaces.
3.2 Continuous Version
Skill discovery. We model the transition kernel as an energy-based model (EBM) [8]:
[…]
where $\psi(s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$ and $\nu(s^{\prime}):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$. The goal is to learn $\psi(s,a)$ and $\nu(s^{\prime})$. $Z(s,a)$ is an intractable partition function. This parametrization provides smooth, normalized probabilities with stable gradients, enabling more effective and generalizable representation learning than the unnormalized inner-product model (Eq. (3)).
Proposition 1 (Connection to Skill Definition).
Given the energy-based transition model in Eq. (7), the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. A contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we sidestep the intractable computation of the partition function $Z(s,a)$ in Eq. (7) by employing NCE (Eq. (2)) to learn the representations.
$\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (8)
$s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples.
Connection to behavior dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. (7) offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. (7) results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. (3) is theoretically valid in continuous domains, we adopt the parameterization in Eq. (7) for continuous state and action spaces, as it more effectively supports the learning of skill representations.
Skill-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. (8), we could theoretically get skill sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. A). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a lightweight neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. (3), we obtain $Q(s,a)=f\bigl{(}\psi(s,a)\bigr{)}^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. (6) becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
$\min_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{\exp(f(\psi(s,a))^{\top}u)}{\exp(f(\psi(s,a))^{\top}u)+\Sigma_{i=1}^{k}\exp(f(\psi(s,a^{\prime}_{i}))^{\top}u)}\right],$ (9)
where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
3.3 Understanding Animal Behavior via Skill and Policy Learning
In this section, we discuss how motor skills $\phi$ and skill weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u(t)$ to vary across tasks or time points. These time-varying coefficients dynamically modulate the influence of each skill. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
Scenario (1): Consider a mouse navigating a maze [23], where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between different reward functions to guide its navigation, with the timing of each reward condition known [24]. We first learn shared skills $\phi(s,a)$ using Eq. (3), then fit task-specific policies $\pi(a|s,t)$ using Eq. (5), where each task $t$ corresponds to one of $T$ reward conditions. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common skill representation across tasks.
Scenario (2): A representative case is a freely behaving mouse [1, 2], where the state is defined by pose keypoints and the action by state change—both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. (8). Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. (9), yielding the motor skill $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying skill compositions, our framework allows multiple skills to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl{(}\phi(s,a)^{\top}u(t)\bigr{)}$, each action is a linear combination of basis skills weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—executed concurrently by different body parts. Unlike discrete switching-state models, which assign one behavior per state, our continuous skill-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
3.4 Reward Recovery
After estimating $u(t)$ as the skill weights, we can further infer $w(t)$ in the linearized reward model $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) [25], where both the policy and underlying reward are inferred. In the context of animal behavior [24, 26, 27], identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is finite.
4 Related work
Skill learning from human behavior. Recent work on skill learning for human behavior understanding has primarily focused on robotics, using deep RL and imitation learning to extract reusable primitives from demonstrations [28, 29, 30, 31]. Parallel advances in IRL aim to improve policy inference using low-dimensional bases [32]. However, these approaches are designed for artificial agents in structured settings and do not capture the fine-grained, time-varying structure of naturalistic behavior. In contrast, our work uses RL-based representation learning on real animal data to uncover task-agnostic motor skills and learn a skill-based policy that models continuous, compositional behavior dynamics—capabilities absent in prior work.
Motor primitives learning from animal behavior. Most existing methods for understanding animal behavior identify motor primitives using classification [33, 34], clustering [4, 5], or behavior segmentation techniques [1, 2, 3]. Existing methods label each time point with discrete syllables (e.g., walking, running) and analyze neural correlates of syllable transitions or structure, but treat behavior as rigid and predefined. They fail to capture its compositional, continuously time-varying nature. To our knowledge, no prior work models motor primitives from this flexible perspective. Our approach introduces representation and imitation learning into animal behavior research, enabling dynamic and interpretable modeling of behavior and internal motivation.
5 Experiments
5.1 Application to simulated data in a multi-task gridworld
We evaluate our algorithm on a multi-task gridworld environment. The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$.
To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
As described in Sec. 3.1, we then learn a set of latent skills and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. We assume a total of 64 skills and learn the 64 $\phi$ vectors. Visualizations of the raw features for these skills are shown in the Appendix. B. Using the learned skills, we recover the policy and further infer the reward function used by the agent, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned skills are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned skills and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC skills and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the skill space.
The PC skills exhibit interpretable patterns. For instance, in skill 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This skill corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that skill 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, skill 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other skills and tasks.
Overall, in this gridworld example, we successfully recover reward functions from behavior trajectories, and importantly, the learned $\phi$ and $w$ are both interpretable and aligned with task goals.
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC skills from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC skills to each task. In the PC skill plot, red indicates positive feature values, blue indicates negative.
5.2 Application to animal navigation behavior
5.2.1 Dataset and model setup
We next evaluate our algorithm on real animal behavior data from [23]. In this experiment, a thirsty mouse is trained to navigate in a maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port located at the end of one branch. Although the mouse’s behavior is primarily driven by water foraging after training, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. The state space has the structure of a binary tree and is defined by the mouse’s location. The actions include moving to its left parent, right parent, left child, and right child. Non-existence of the moving target will end in staying still.
We first apply the EM-based segmentation algorithm from [24] to divide the behavior into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting.
To uncover a task-agnostic set of skills capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. The learned $\phi$ vectors serve as a policy basis, encoding movement tendencies such as approaching or avoiding certain regions of the maze.
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant skills for the water and home tasks respectively. Each skill indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight skills for water seeking, and orange for home seeking.
5.2.2 Results
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. In the home-seeking task, a distinct peak appears at the home cage. In the exploration task, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned skills by applying PCA to obtain PC skills. See the Appendix. C for raw features of skills. Fig. 3B displays two top-contributing skills for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related skills promote movement toward the water port, while the home-related skills guide navigation to the cage. Notably, the skills important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization.
These results show that our model accurately recovers multiple reward functions from real behavior and, importantly, learns interpretable skills whose contributions to each task are both distinct and behaviorally meaningful.
5.3 Application to animal free-moving behavior
5.3.1 Dataset and model setup
In this section, we apply our method to a dataset of free-moving mouse behaviors to extract motor skills and analyze pose dynamics, specifically how complex pose sequences are dynamically composed of underlying motor primitives. The dataset [2] contains the keypoint coordinates of eight mouse body parts, including four points along the spine, the head, the nose, and both ears. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of the state, computed as $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP.
As outlined in Sec. 3.2, we parameterize the transition function as $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})\right)$, and use NCE to learn the feature functions $\psi(s,a)$ and $\nu(s^{\prime})$. We then apply a transformation function $f$ and again use NCE to learn the skill representation $\phi(s,a)=f(\psi(s,a))$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. We optimize $f$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned skill representation $\phi$ and the temporal weights $u(t)$.
We split the full dataset into training and test trajectories in an 8:2 ratio. We first learn both $f$ and $u(t)$ on the training set. Then, given the learned $f$, we estimate $u(t)$ on the test set. Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
We compare our SKIL method with a baseline model, Keypoint-MoSeq [2], a state-of-the-art behavior segmentation algorithm that partitions continuous pose trajectories into discrete behavioral syllables. For a fair comparison, we train Keypoint-MoSeq on the training set to learn its parameters and then use it to infer behavioral syllables, i.e., perform segmentation, on the test trajectories.
5.3.2 Results
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. We use the log-likelihood of the autoregressive model in Keypoint-MoSeq as its prediction score and the negative energy function $f(\psi(s,a))^{\top}u(t)$ as our prediction score, respectively.
We repeat the experiment 10 times and report the results as a box plot in Fig. 4A. Our model achieves higher AUC scores on both training and test sets (paired t-test, $p_{train}=0.0037,p_{test}=0.0046$), demonstrating a stronger ability to distinguish positive from negative samples. This suggests that SKIL more accurately captures time-varying pose dynamics through smoothly evolving motor primitives, while Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity.
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key skills $\phi(s,a)$ associated with example pose dynamics. We examine the time-varying weights $u(t)$ and compare them with the behavioral syllable segmentation produced by Keypoint-MoSeq (Fig. 4B). We present five example clips of observed animal behavior, each consisting of five consecutive frames. Each clip is characterized by a unique combination of motor skills. For each clip, we show the top 1-2 most dominant skills. For each skill, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that skill. The movement semantics of each skill are labeled above the visualizations. A more comprehensive visualization of all learned skills is included in the Appendix. D.
By combining the discovered skills with real animal behavior, we assess the interpretability of each skill in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward skills (skill 1 and skill 2). A stronger movement at the head is reflected by a higher value of skill 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of skill 3 and skill 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of skill 5 and skill 6. Across all skills, skill 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned skills and their temporal weights $u(t)$.
However, when we examine the segmentation produced by Keypoint-MoSeq and the syllable dynamics (Fig. 4C), we find inconsistencies. For example, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with similar approaches, SKIL provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor skills. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor primitives.
Figure 4: A. AUC on training and test sets. B. We present example learned skills, behavior clips, $u(t)$, and segmentation from Keypoint-MoSeq. C. Syllables identified by Keypoint-MoSeq (adopted from [2]).
6 Discussion
We introduce the SKIL framework that advances animal behavior modeling by offering a flexible, interpretable alternative to traditional discrete behavioral segmentation methods. It learns task-agnostic motor primitives (skill sets) that dynamically compose into complex behaviors. SKIL effectively recovers meaningful skills, reconstructs reward functions, and captures nuanced behavioral dynamics, providing a principled and quantitative model of how complex policies arise from fundamental motor components.
Several limitations remain to be addressed in future work. First, the accuracy of inferred skills is sensitive to input data quality, as occlusions or tracking errors can degrade performance. Additionally, training the energy-based models can also be computationally intensive, potentially restricting real-time or large-scale applications. Lastly, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "skills" and specific neural circuit dynamics still requires further experimental validation.
Beyond advancing animal behavior research, SKIL has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Appendix A Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
By simple algebra, we obtain the quadratic potential function,
$P(s^{\prime}|s,a)$ $\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (10)
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature [6, 35] and obtain the spectral decomposition of Eq. (7) as
$P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (11)
where $\omega\sim\mathcal{N}(0,I)$ is the frequency in the Fourier domain, and
$\phi_{\omega}(s,a)$ $=\exp\left(-i\omega^{\top}\psi(s,a)\right)\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (12)
$\mu_{\omega}(s^{\prime})$ $=\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (13)
Note that Eq. (11) needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
$P(s^{\prime}|s,a)\approx\frac{1}{M}\sum_{i=1}^{M}\phi_{\omega_{i}}(s,a)\mu_{\omega_{i}}(s^{\prime})q(s^{\prime}).$ (14)
Introduce vectors $\phi(s,a)$ and $\mu(s^{\prime})$ such that
$\phi(s,a)$ $:=\frac{1}{\sqrt{M}}\left[\phi_{\omega_{1}}(s,a),\phi_{\omega_{2}}(s,a),...,\phi_{\omega_{M}}(s,a)\right],$ (15)
$\mu(s^{\prime})$ $:=\frac{1}{\sqrt{M}}\left[\mu_{\omega_{1}}(s^{\prime}),\mu_{\omega_{2}}(s^{\prime}),...,\mu_{\omega_{M}}(s^{\prime})\right].$ (16)
Then it’s straightforward to see that,
$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))\approx P(s^{\prime}|s,a).$ (17)
Hence, Eq. (7) can, in principle, yield the skill representation introduced earlier.
Appendix B Learned Skills in the Multi-task Gridworld Dataset
Figure 5: A. State-action maps for all 64 skills. B. Reward weight $w$ for all 64 skills.
We visualize all original 64 skills (Fig. 5) introduced in Sec. 5.1. It shows some meaningful patterns as mentioned before. For example, skill 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this skill and used skill 39 negatively instead. The complex many-to-many relationship between skills and tasks informs us of the redundancy in the original skills, which inspires us to use PCA to analyze the principal components of the skill space and simplify the skill weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the skill space.
Appendix C Learned Skills in the Animal Navigation Behavior Dataset
Figure 6: A. Left. The value for each $(s,a)$ pair in skill 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in skill 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
In the original skills of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the skill. Given the redundancy of the skill sets, as in Appendix. B, we perform PCA to analyze the principle components of the skill space and simplify the skill representation. To show the effect of PCA, we plot one skill (skill 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in skill 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the skill becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all skills and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each skill is called the “effective dimension.” The effective dimension is calculated across all skills (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Appendix D Learned Skills in the Animal Free-moving Behavior Dataset
Figure 7: A. Policy weight $u(t)$ for all 64 skills. B. The motion field for all 64 skills learned from the video, computed by averaging the states and actions that most strongly activate each skill.
We show all skills learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec. 5.3.2. We have completed the visualization of those skills that were previously omitted due to their perceived lack of importance. Due to the increased number of displayed skills, we had to renumber each skill. We show the present skill number above the skill motion field figure, and previous numbers (if applicable) in the parentheses following the present number.

### v2

@@
-Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior
+Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior
 Abstract
-Animals flexibly recombine a finite set of core motor primitives to meet diverse task demands, but existing behavior-segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions.
-To reflect the animal behavior generation procedure, we introduce skill-based imitation learning (SKIL) for behavior understanding, a reinforcement learning-based imitation framework that (1) infers interpretable skill sets, i.e., latent basis functions of behavior, by leveraging representation learning on transition probabilities, and (2) parameterizes policies as dynamic mixtures of these skills.
-We validate our approach on a simple grid world, a discrete labyrinth, and unconstrained videos of freely moving animals.
-Across tasks, it identifies reusable skill components, learns continuously evolving compositional policies, and generates realistic trajectories beyond the capabilities of traditional discrete models.
-By exploiting generative behavior modeling with compositional representations, our method offers a concise, principled account of how complex animal behaviors emerge from dynamic combinations of fundamental motor primitives.
+Animals flexibly recombine a finite set of core motor motifs to meet diverse task demands, but existing behavior segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions.
+To better capture the continuous structure of behavior generation, we introduce motif-based continuous dynamics (MCD) discovery, a framework that (1) uncovers interpretable motif sets as latent basis functions of behavior by leveraging representations of behavioral transition structure, and (2) models behavioral dynamics as continuously evolving mixtures of these motifs.
+We validate MCD on a multi-task gridworld, a labyrinth navigation task, and freely moving animal behavior.
+Across settings, it identifies reusable motif components, captures continuous compositional dynamics, and generates realistic trajectories beyond the capabilities of traditional discrete segmentation models.
+By providing a generative account of how complex animal behaviors emerge from dynamic combinations of fundamental motor motifs, our approach advances the quantitative study of natural behavior.
 1 Introduction
-Animal behavior is composed of recurring patterns that reflect underlying motor primitives like back grooming, running, and sniffing.
-Behavior segmentation methods [1, 2, 3, 4, 5] seek to uncover structured patterns in behavior by dividing continuous pose trajectories into discrete units known as stereotyped behavioral syllables.
-These segmented units can then be used to build structured representations of movement for downstream neurobehavioral analysis.
-However, existing behavior segmentation methods face three major limitations.
+A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings.
+Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu & Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables.
+Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu & Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) switching-dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022).
+The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
+However, existing behavior segmentation methods face four major limitations.
 First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions.
-Second, they often abstract actions into high-level syllables, overlooking shared motion components across body parts.
+Second, they often extract actions as abstract syllables that fail to capture how individual body parts contribute to different motions.
 For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns.
 Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior.
-Third, most segmentation models are either non-generative (e.g., clustering [4, 5]) or rely on restrictive generative assumptions (e.g., linear dynamics and hidden Markov model [1, 2, 3]), often leading to unrealistic synthesized behaviors.
-To address these limitations, we introduce a new perspective: modeling behavior as emerging from an agent following a reinforcement learning (RL) policy.
-We study behavioral dynamics and motor primitives by inferring the animal’s policy through an RL–based imitation learning framework.
-It not only enables more realistic behavior generation through RL but also allows us to discover motor primitives as reusable skill sets that drive behavior toward internal goals.
-By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor primitives that go beyond the constraints of discrete segmentation.
-Concretely, we first learn interpretable latent representations from behavior trajectories, which we refer to as skill sets.
-These represent low-level motion patterns that act as modular building blocks of behavior.
-Such skills can involve different body parts.
-By flexibly combining these skill sets, animals can generate a wide range of complex, high-level behaviors.
-For example, face grooming (forepaw-to-face movements) and body grooming (repetitive strokes along the torso) share common grooming-related skills, while also engaging distinct body parts specific to each movement.
-We leverage existing advances in representation learning in RL [6, 7, 8] to learn skill sets from transition probability functions.
-Then, we parametrize policies using these skill sets and apply imitation learning [9] to train the policies from animal behavior demonstrations.
-Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor skills.
-Our method is referred to as SKIL, short for Skill-Based Imitation Learning.
-The fine-grained skill representation enables highly flexible behavioral characterization through the policy.
-Critically, over long behavioral trajectories, we assume an animal draws from a fixed set of core motor skills, regardless of the specific movements performed [10, 11].
-The policy captures how the contribution of each skill evolves continuously over time, reflecting dynamic behavioral changes.
-Due to the spatial specificity of different motor components, some skills may be engaged briefly, while others persist over longer periods.
-Moreover, multiple skills can be active simultaneously to construct ongoing movement (Fig.
+Third, most segmentation models are either non-generative (e.g., clustering (Hsu & Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors.
+Fourth and most importantly, classification/clustering-based methods ignore temporal dependencies and fail to capture motifs reflecting true pose dynamics, while dynamics-based methods impose linear or nonlinear dynamics assumptions that may not match actual animal behavior, making such assumptions potentially harmful.
+To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework.
+We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework.
+It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards.
+By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation.
+Note that Aldarondo et al.
+(2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
+We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash & Hochner, 2005).
+Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics.
+Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2014; Ren et al., 2023; Shribak et al., 2024).
+These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts.
+For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts.
+These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors.
+Finally, we apply imitation learning to train motif-based policies from demonstrations.
+This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics.
+As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
+Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor motifs.
+Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes.
+Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig.
 1).
-This flexible, compositional view of behavior cannot be achieved using traditional discrete behavioral syllables.
-SKIL thus offers a more nuanced understanding of how complex actions are generated from dynamic combinations of fundamental motor primitives.
-Such insight enables researchers to test whether specific fine-grained motor primitives depend on certain local neural circuits, offering a more precise understanding of movement control and its underlying neural mechanisms.
-In the remainder of the paper, we introduce preliminaries in Sec.
-2, which pave the way for our methods for animal behavior understanding in Sec.
-3.
-We review related work in Sec.
-4 and demonstrate the advantage of our methods empirically in Sec.
-5.
-Finally, we conclude with a discussion in Sec.
-6.
+This flexible, compositional view cannot be achieved with traditional discrete syllables.
+MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits.
+Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches.
+Note that there is a rich literature in robotics on learning skills through imitation learning (Paraschos et al., 2013; Lioutikov et al., 2017; Li et al., 2017; Ajay et al., 2021; Peng et al., 2022; Kuang et al., 2025).
+We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper.
+Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022).
+To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
+  • We introduce the first RL-based IL framework for behavior segmentation, a fundamental advance since RL naturally treats behavior as a decision-making process shaped by policies and rewards.
+Unlike dynamics-based methods, it explains why behaviors occur, not just how they unfold.
+  • Within this framework, we propose RL-based representation learning to discover motif-based policies.
+The learned motifs and policies do not rely on dynamics assumptions or any model assumptions.
+They can faithfully characterize behavioral dynamics without the mismatch issues of prior segmentation methods.
+  • Our method reveals the continuous nature of animal behavior, moving beyond the discrete segmentation assumptions of existing methods.
+  • It provides a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
 2 Preliminaries
 Markov Decision Processes (MDP).
-To define skills that characterize mouse behavior, we begin by modeling the observed behavioral trajectories within the framework of MDPs.
-Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and the animal’s condition—for example, as defined by pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon.
+To define motifs that characterize animal behaviors, we begin by modeling the observed behavioral trajectories within the framework of MDPs.
+Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and an animal’s condition—for example, positions of pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon.
 A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step.
-We assume that the animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior.
+We assume that an animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior.
 The behavior is governed by a policy that seeks to maximize this internal reward.
@@
 $Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
-Imitation Learning.
-The goal of imitation learning(IL) is to find a policy $\pi$ that mimics the behavior of the animal demonstrations.
+Offline Imitation Learning.
+We use imitation learning(IL) to find a policy $\pi$ that mimics animal behavior.
 In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
-[…]
-$f^{*}=\mathop{\arg\max}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\log\frac{f(x_{i},c_{i})}{f(x_{i},c_{i})+\sum_{j=1}^{k}f(y_{i,j},c_{i})},$ (2)
-where $\{(x_{i},c_{i})\}_{i=1}^{n}$ are observed samples, $\{y_{i,j}\}_{i\in[n],j\in[k]}$ are artificially generated data sampled from a predefined noise distribution $q(y)$.
-Here, $k$ represents the ratio of negative to positive samples.
-This objective can be interpreted as a multi-class extension for distinguishing real data from noise.
-3 Skill-based Imitation learning (SKIL) for Behavior Understanding
-Given the MDP definition, we frame skill discovery from a control-theoretic perspective.
-In this view, skills are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
-Definition 1 (Skill Set).
-Given an arbitrary transition kernel $P(\cdot|s,a)$ in a linear (or low-rank) MDP [15, 16, 17], we can express it via a spectral decomposition:
+3 Motif-based Continuous Dynamics (MCD) Discovery
+Given the MDP definition, we frame motif discovery from a control-theoretic perspective.
+In this view, motifs are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
+Definition 1 (Motif Set).
+Given an arbitrary transition kernel $P(\cdot|s,a)$ in an MDP, we can express it via a spectral decomposition:
 $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime}),$ (2)
-where $\phi:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:{\mathcal{S}}\rightarrow\mathbb{R}^{d}$, and $q(s^{\prime})$ is some parametrized probability measure supported on the state space $\mathcal{S}$.
-We define the function $\phi$ as the skill set.
+where $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:\mathcal{S}\rightarrow\mathbb{R}^{d}$, and $q\in\Delta(\mathcal{S})$ is a parametrized probability distribution over the state space.
+We define the function $\phi$ as the motif set.
 The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
-Because animal behavior (e.g., navigation, pose dynamics) follows structured, stereotyped patterns that evolve on low-dimensional manifolds, it can be effectively approximated by low-rank MDPs even in the presence of underlying nonlinearities.
-Given the skill definition, substituting Eq. (3) and the linearized reward model into the Bellman equation (Eq. (1)), we get:
-$\!Q(s,a)$ $=r(s,a)+\gamma\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
-$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (4)
+Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2023; Zhang et al., 2022; Shribak et al., 2024).
+We adopt this approach here to define motifs given the transition kernel and define rewards accordingly.
+Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
+Given the motif definition, substituting Eq. 2 and the linearized reward model into the Bellman equation (Eq. 1), we get:
+$Q(s,a)=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (3)
 where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$.
-Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of skill features $\phi(s,a)$, offering a convenient way to link skills to the policy.
-Following the maximum entropy reinforcement learning framework [18], we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy.
-Under this assumption, the optimal policy $\pi(a|s)$ can be shown to follow:
+Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of motif features $\phi(s,a)$, offering a convenient way to link motifs to the policy.
+Following the maximum entropy reinforcement learning framework (Haarnoja et al., 2018), we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy.
+Under this assumption, the optimal max-entropy policy $\pi(a|s)$ can be shown to follow:
 $\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
 where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$ is the entropy.
-Thereby, the skill function $\phi(s,a)$ not only spans the space of $Q$-functions but also provides a universal basis for constructing the policy $\pi(a|s)$ via Eq.
-(5).
-Since $\phi(s,a)$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward function or task.
-In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor skills available to the animal, while the weight vector $u$ captures task-specific modulations required to produce behavior consistent with different goals.
-With the skill set established as a foundational representation for both value functions and policies, we are equipped to interpret behavioral trajectories through the lens of skill alignment.
-Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (skill discovery), as well as $u$ (skill-based policy learning) from state-action-next-state observations.
-The learning procedure differs depending on the nature of the behavior data.
+Proposition 1.
+The policy in Eq.
+4 is not based on any model assumption but emerges naturally as the max-entropy policy based on spectral decomposition of the transition kernel.
+Furthermore, the learned motifs $\phi$ can represent any max-entropy policy through an appropriate choice of $u$.
+The reason we define $\phi(s,a)$ as motifs is that $\phi$ provides the linear basis for the environment transition, policies, and rewards.
+Policies characterize behavioral dynamics by describing action tendencies conditioned on state (how behaviors evolve).
+Combined with the environment dynamics $P(s^{\prime}|s,a)$, it induces the transition distribution $P(s^{\prime}|s)=\sum_{a}P(s^{\prime}|s,a)\,\pi(a|s)$, which governs the evolution of behavior trajectories, as is often directly modeled in dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024).
+Rewards, in turn, reflect the underlying driving factors of these trajectories (why behaviors evolve).
+Moreover, because $\phi$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward or task.
+In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor motifs available to animals, while the weight vector $u$ captures task-specific modulations required to produce behavior aligned with different goals.
+Thus, we can interpret behavioral trajectories through the lens of motifs $\phi$.
+From Def.
+1 and Prop.
+1, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state.
+Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
+Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning that characterizes the continuous behavioral dynamics) from demonstrations.
+The learning procedure differs depending on the nature of the behavior data (i.e.
+discrete or continuous).
 3.1 Discrete Version
-Skill discovery.
-For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) [19, 20, 21] or spectral decomposition representation [7, 22] to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$.
-The resulting skill set $\phi(s,a)$ is then used in the subsequent policy learning stage.
-Skill-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. (5), using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
+Motif discovery.
+For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) (Golub & Reinsch, 1971; Golub & Van Loan, 2013; Trefethen & Bau, 2022) or spectral decomposition representation (Ren et al., 2023; HaoChen et al., 2021) to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$.
+The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
+Motif-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. 4, using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
 $\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
-Limitations for continuous data.
+3.2 Continuous Version
 While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons.
-First, in the skill discovery step (Eq.
-(3)), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals [2].
+First, in the motif discovery step (Eq.
+2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024).
 For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise.
 This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable.
-Second, in skill-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq.
-(6) is easy to compute.
+Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq.
+5 is easy to compute.
 But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate.
-In light of these challenges, we adopt an alternative approach to learn the skill representations and policy in continuous state-action spaces.
-3.2 Continuous Version
-Skill discovery. We model the transition kernel as an energy-based model (EBM) [8]:
-[…]
-where $\psi(s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$ and $\nu(s^{\prime}):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$.
-The goal is to learn $\psi(s,a)$ and $\nu(s^{\prime})$.
-$Z(s,a)$ is an intractable partition function.
-This parametrization provides smooth, normalized probabilities with stable gradients, enabling more effective and generalizable representation learning than the unnormalized inner-product model (Eq.
-(3)).
-Proposition 1 (Connection to Skill Definition).
-Given the energy-based transition model in Eq.
-(7), the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
-Appendix. A contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
-To learn $\psi$ and $\nu$, we sidestep the intractable computation of the partition function $Z(s,a)$ in Eq. (7) by employing NCE (Eq. (2)) to learn the representations.
+In light of these challenges, we adopt an alternative approach to learn the motif representations and policy in continuous state-action spaces.
+Motif discovery. We model $P(s^{\prime}|s,a)$ as an energy-based model (EBM) (Shribak et al., 2024):
+$P(s^{\prime}|s,a)=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),Z(s,a)=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
+where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{g}$ and $\nu:\mathcal{S}\to\mathbb{R}^{g}$ are neural-network feature maps.
+Here, $Z(s,a)$ is an intractable partition function.
+Compared to the unnormalized inner-product model (Eq.
+2), this EBM formulation yields smooth, normalized probabilities and stable gradients, leading to more effective and generalizable motif representations.
+Proposition 2 (Connection to Motif Definition).
+Given the EBM model in Eq.
+6, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
+Appendix. C contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
+To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Ma & Collins, 2018; Gutmann & Hyvärinen, 2010; 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function.
+In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq.
+6 by solving
 $\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (7)
-$s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples.
-Connection to behavior dynamics.
-With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$.
+where $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from distribution $\rho$.
+Connection to behavioral dynamics.
+With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq.
+6.
 By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier.
 Thus, Eq.
-(7) offers a more general framework that extends traditional dynamics models for studying behavior.
+6 offers a more general framework that extends traditional dynamics models for studying behavior.
 Moreover, in practice, using Eq.
-(7) results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data.
+6 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data.
 Thus, while Eq.
-(3) is theoretically valid in continuous domains, we adopt the parameterization in Eq.
-(7) for continuous state and action spaces, as it more effectively supports the learning of skill representations.
-Skill-based policy learning.
+2 is theoretically valid in continuous domains, we adopt the parameterization in Eq.
+6 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
+Motif-based policy learning.
 After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq.
-(8), we could theoretically get skill sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix.
-A).
+7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix.
+C).
 However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly.
-Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a lightweight neural network, and learn it via policy learning.
+Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning.
 Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data.
 By applying $\phi=f(\psi)$ to Eq.
-(3), we obtain $Q(s,a)=f\bigl{(}\psi(s,a)\bigr{)}^{\top}u$.
+3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
 As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq.
-(6) becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space.
+5 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space.
 This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$.
@@
 where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
-3.3 Understanding Animal Behavior via Skill and Policy Learning
-In this section, we discuss how motor skills $\phi$ and skill weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u(t)$ to vary across tasks or time points.
-These time-varying coefficients dynamically modulate the influence of each skill.
+3.3 Understanding Animal Behavior via Motif and Policy Learning
+In this section, we discuss how motor motifs $\phi$ and motif weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u$ to vary across tasks $t$ or time points $t$.
+As the motif coefficients, $u(t)$ would dynamically modulate the influence of each motif on the final policy.
 We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
-Scenario (1): Consider a mouse navigating a maze [23], where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay).
-We assume the animal switches between different reward functions to guide its navigation, with the timing of each reward condition known [24].
-We first learn shared skills $\phi(s,a)$ using Eq.
-(3), then fit task-specific policies $\pi(a|s,t)$ using Eq.
-(5), where each task $t$ corresponds to one of $T$ reward conditions.
-This yields $T$ sets of weights $u(t)$, one per task, while sharing a common skill representation across tasks.
-Scenario (2): A representative case is a freely behaving mouse [1, 2], where the state is defined by pose keypoints and the action by state change—both continuous.
+The concrete results will be later shown in Sec.
+4.2 and Sec.
+4.3 respectively.
+In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task or time point.
+Scenario (1): Consider a mouse navigating a maze (Rosenberg et al., 2021), where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay).
+We assume the animal switches between $T$ strategies, each associated with a unique reward, that guide its navigation, with the timing of each reward condition known from Ke et al.
+(2025).
+We first learn shared motifs $\phi(s,a)$ using Eq.
+2, then fit task-specific policies $\pi(a|s,t)$ using Eq.
+5, where each task $t$ corresponds to one of $T$ strategies.
+This yields $T$ sets of weights $u(t)$, one per task, while sharing a common motif set across tasks.
+Scenario (2): A representative case is a freely behaving mouse (Wiltschko et al., 2015; Weinreb et al., 2024), where the state is defined by pose keypoints and the action by state change–both continuous.
 We first learn $\psi(s,a)$ from pose trajectories using Eq.
-(8).
+7.
 Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq.
-(9), yielding the motor skill $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$.
+8, yielding the motor motif $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$.
 Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
-Beyond capturing smoothly time-varying skill compositions, our framework allows multiple skills to be active simultaneously.
-Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl{(}\phi(s,a)^{\top}u(t)\bigr{)}$, each action is a linear combination of basis skills weighted by $u(t)$, enabling overlapping and composable behaviors.
-For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—executed concurrently by different body parts.
-Unlike discrete switching-state models, which assign one behavior per state, our continuous skill-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
+Beyond capturing smoothly time-varying motif compositions, our framework allows multiple motifs to be active simultaneously.
+Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl(\phi(s,a)^{\top}u(t)\bigr)$, each action is a generalized linear combination of basis motifs weighted by $u(t)$, enabling overlapping and composable behaviors.
+For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—recruiting different motifs concurrently.
+Unlike discrete switching-state models, which assign one behavior per state, our continuous motif-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
+In Sec.
+4, we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
 3.4 Reward Recovery
-After estimating $u(t)$ as the skill weights, we can further infer $w(t)$ in the linearized reward model $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$.
+After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$.
 This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals.
-Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) [25], where both the policy and underlying reward are inferred.
-In the context of animal behavior [24, 26, 27], identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes.
-Since $V(s,t)$ can be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is finite.
-4 Related work
-Skill learning from human behavior.
-Recent work on skill learning for human behavior understanding has primarily focused on robotics, using deep RL and imitation learning to extract reusable primitives from demonstrations [28, 29, 30, 31].
-Parallel advances in IRL aim to improve policy inference using low-dimensional bases [32].
-However, these approaches are designed for artificial agents in structured settings and do not capture the fine-grained, time-varying structure of naturalistic behavior.
-In contrast, our work uses RL-based representation learning on real animal data to uncover task-agnostic motor skills and learn a skill-based policy that models continuous, compositional behavior dynamics—capabilities absent in prior work.
-Motor primitives learning from animal behavior.
-Most existing methods for understanding animal behavior identify motor primitives using classification [33, 34], clustering [4, 5], or behavior segmentation techniques [1, 2, 3].
-Existing methods label each time point with discrete syllables (e.g., walking, running) and analyze neural correlates of syllable transitions or structure, but treat behavior as rigid and predefined.
-They fail to capture its compositional, continuously time-varying nature.
-To our knowledge, no prior work models motor primitives from this flexible perspective.
-Our approach introduces representation and imitation learning into animal behavior research, enabling dynamic and interpretable modeling of behavior and internal motivation.
+Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008), where both the policy and underlying reward are inferred from demonstrations.
+In the context of animal behavior (Ke et al., 2025; Zhu et al., 2024; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes.
+Since $V(s,t)$ can only be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is discrete and finite.
 4 Experiments
 4.1 Application to simulated data in a multi-task gridworld
-We evaluate our algorithm on a multi-task gridworld environment.
 The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig.
@@
 In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$.
-To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
+See Appendix.
+D for more details for data generation.
 As described in Sec.
-3.1, we then learn a set of latent skills and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$.
-We assume a total of 64 skills and learn the 64 $\phi$ vectors.
-Visualizations of the raw features for these skills are shown in the Appendix.
-B.
-Using the learned skills, we recover the policy and further infer the reward function used by the agent, as described in Sec.
+3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$.
+Because the computational complexity only scales linearly with the number of motifs, we assume a total of 64 motifs and learn the 64 $\phi$ vectors to cover the motif space as much as possible.
+Visualizations of these motifs are shown in the Appendix.
+D.
+Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec.
 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
 This recovered reward (Fig.
-2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned skills are sufficient for accurately reconstructing the reward function from behavior data.
-To better interpret the learned skills and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig.
+2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
+To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig.
 2B).
-Therefore, we visualize the top 8 PC skills and their corresponding task-specific coefficients $w(t)$ in Fig.
-2D, treating them as basis vectors spanning the skill space.
-The PC skills exhibit interpretable patterns.
-For instance, in skill 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values.
-This skill corresponds to moving away from the middle-right grid toward the bottom-left.
-Examining the $w$ matrix, we see that skill 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task.
-In contrast, skill 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal.
-Similar interpretations can be made for other skills and tasks.
-Overall, in this gridworld example, we successfully recover reward functions from behavior trajectories, and importantly, the learned $\phi$ and $w$ are both interpretable and aligned with task goals.
+Therefore, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig.
+2D, treating them as basis vectors spanning the motif space.
+The PC motifs exhibit interpretable patterns.
+For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values.
+This motif corresponds to moving away from the middle-right grid toward the bottom-left.
+Examining the $w$ matrix, we see that motif 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task.
+In contrast, motif 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal.
+Similar interpretations can be made for other motifs and tasks.
+In this gridworld experiment, we successfully recover reward functions from behavior trajectories, and, importantly, the learned $\phi$ and $w$ are effectively deployed in different task settings.
 Figure 2: A.
@@
 D.
-Reward weight $w$ and top 8 PC skills from the $\phi$ matrix.
-Reward weights indicate the contribution of the top 8 PC skills to each task.
-In the PC skill plot, red indicates positive feature values, blue indicates negative.
+Reward weight $w$ and top 8 PC motifs from the $\phi$ matrix.
+Reward weights indicate the contribution of the top 8 PC motifs to each task.
+In the PC motif plot, red indicates positive feature values, blue indicates negative.
 4.2 Application to animal navigation behavior
-5.2.1 Dataset and model setup
-We next evaluate our algorithm on real animal behavior data from [23].
-In this experiment, a thirsty mouse is trained to navigate in a maze (Fig.
-3A), starting each trial from the central home cage and attempting to reach a water port located at the end of one branch.
-Although the mouse’s behavior is primarily driven by water foraging after training, it also exhibits exploration of unvisited areas and returns to the home cage for shelter.
+Dataset and model setup.
+We next evaluate our method on a real animal behavior dataset from Rosenberg et al.
+(2021).
+In this experiment, a thirsty mouse is trained to navigate in a binary-tree maze (Fig.
+3A), starting each trial from the central home cage and attempting to reach a water port at one leaf node.
+The state space is defined by the mouse’s location on the tree.
+The actions include moving to its left parent, right parent, left child, and right child.
+Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter.
 This complex behavior cannot be captured by a single reward function.
-The state space has the structure of a binary tree and is defined by the mouse’s location.
-The actions include moving to its left parent, right parent, left child, and right child.
-Non-existence of the moving target will end in staying still.
-We first apply the EM-based segmentation algorithm from [24] to divide the behavior into three interpretable tasks: water seeking, home seeking, and exploration.
+Studying this dataset allows us to discover motifs shared across multiple reward functions and policies.
+This in turn tests whether complex navigation behavior, under multiple competing motivations, can be distilled into a small set of reusable motifs that provide interpretable insight into animal decision-making.
+We first apply the segmentation algorithm from Ke et al.
+(2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration.
 The algorithm could also infer reward maps for these tasks, which are shown in Fig.
@@
 Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting.
-To uncover a task-agnostic set of skills capable of constructing all three reward functions, we apply our model as described in Sec.
+To uncover a task-agnostic set of motifs capable of constructing all three reward functions, we apply our model as described in Sec.
 3.1.
-The learned $\phi$ vectors serve as a policy basis, encoding movement tendencies such as approaching or avoiding certain regions of the maze.
+Again, to cover the motif space as much as possible, we set the number of motifs $d=128$.
 Figure 3: A.
@@
 B.
-Two dominant skills for the water and home tasks respectively.
-Each skill indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value.
+Two dominant motifs for the water and home tasks respectively.
+Each motif indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value.
 C.
-$w$ values for all tasks, colored by task; blue stars highlight high-weight skills for water seeking, and orange for home seeking.
-5.2.2 Results
+$w$ values for all tasks, colored by task; blue stars highlight high-weight motifs for water seeking, and orange for home seeking, all shown in (B).
+Results.
 To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
@@
 In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it.
-In the home-seeking task, a distinct peak appears at the home cage.
-In the exploration task, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
-We further visualize the learned skills by applying PCA to obtain PC skills.
+During home-seeking, a distinct peak appears at the home cage.
+When exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
+We further visualize the learned motifs by applying PCA to obtain PC motifs.
 See the Appendix.
-C for raw features of skills.
+E for raw features of motifs.
 Fig.
-3B displays two top-contributing skills for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig.
+3B displays two top-contributing motifs for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig.
 3C).
-The water-related skills promote movement toward the water port, while the home-related skills guide navigation to the cage.
-Notably, the skills important for one task have minimal or negative contributions to the other (Fig.
+The water-related motifs promote movement toward the water port, while the home-related motifs guide navigation to the cage.
+Notably, the motifs important for one task have minimal or negative contributions to the other (Fig.
 3C), indicating clear functional specialization.
-These results show that our model accurately recovers multiple reward functions from real behavior and, importantly, learns interpretable skills whose contributions to each task are both distinct and behaviorally meaningful.
+Fig.
+3C also shows non-task-specific motifs with similar weights in both water- and home-seeking tasks.
+These results show that our model not only recovers multiple reward functions from real behavior but also learns interpretable motifs whose contributions to each task are distinct and behaviorally meaningful.
+Unlike previous reward discovery on this dataset (Ashwood et al., 2022; Ke et al., 2025), which cannot identify such motifs, our approach reveals how reward maps can be decomposed into smaller, reusable action components.
+This decomposition offers a mechanistic view of how local decision processes combine to produce strategies such as water-seeking or home-seeking.
 4.3 Application to animal free-moving behavior
-5.3.1 Dataset and model setup
-In this section, we apply our method to a dataset of free-moving mouse behaviors to extract motor skills and analyze pose dynamics, specifically how complex pose sequences are dynamically composed of underlying motor primitives.
-The dataset [2] contains the keypoint coordinates of eight mouse body parts, including four points along the spine, the head, the nose, and both ears.
+Dataset and model setup.
+We finally apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024) to extract motor motifs and analyze fine-grained pose dynamics.
+This dataset contains the keypoint coordinates of eight mouse body parts, including the head, the nose, both ears, and four spine nodes.
 Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part.
-The action at each timestep is defined as the velocity of the state, computed as $a_{t}=(s_{t+1}-s_{t})/\delta t$.
+The action at each timestep is defined as the velocity of state, $a_{t}=(s_{t+1}-s_{t})/\delta t$.
 We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset).
 We formulate the data using a continuous MDP.
+Studying this dataset allows us to ask whether free-moving behaviors, which often appear as mixtures of grooming, locomotion, and postural adjustments, can be represented as combinations of a compact set of task-agnostic motor motifs.
 As outlined in Sec.
-3.2, we parameterize the transition function as $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})\right)$, and use NCE to learn the feature functions $\psi(s,a)$ and $\nu(s^{\prime})$.
-We then apply a transformation function $f$ and again use NCE to learn the skill representation $\phi(s,a)=f(\psi(s,a))$ and the time-dependent weights $u(t)$.
+3.2, we use NCE to learn the motif representation $\phi(s,a)$ and the time-dependent weights $u(t)$.
 To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$.
-We optimize $f$ and $u(t)$ using coordinate descent, updating them alternately.
+To perform more stable optimization, we optimize $\phi(s,a)$ and $u(t)$ using coordinate descent, updating them alternately.
 For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses.
-Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned skill representation $\phi$ and the temporal weights $u(t)$.
-We split the full dataset into training and test trajectories in an 8:2 ratio.
-We first learn both $f$ and $u(t)$ on the training set.
-Then, given the learned $f$, we estimate $u(t)$ on the test set.
-Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
-We compare our SKIL method with a baseline model, Keypoint-MoSeq [2], a state-of-the-art behavior segmentation algorithm that partitions continuous pose trajectories into discrete behavioral syllables.
-For a fair comparison, we train Keypoint-MoSeq on the training set to learn its parameters and then use it to infer behavioral syllables, i.e., perform segmentation, on the test trajectories.
-5.3.2 Results
+Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned motif representation $\phi$ and the temporal weights $u(t)$.
+With respect to the choice of the number of motifs, we find in practice that the performance grows more slowly once past $d=64$, so we choose this as an optimal number.
+See Appendix.
+F for more training details.
+We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), as a representative for switching-dynamics-based segmentation methods; and (2) SemiSeg (Whiteway et al., 2021), as a representative for clustering-based behavior segmentation methods.
+Note that Keypoint-MoSeq is regarded as a SOTA approach, because it extends the autoregressive hidden Markov model (AR-HMM) and MoSeq (Wiltschko et al., 2015), and has been shown to outperform B-SOiD (Hsu & Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014).
+We also include OPAL (Ajay et al., 2021), a representative autoencoder-based motif learning algorithm from robotics, as a baseline to highlight the advantages of our method and why robotics approaches are ill-suited for behavioral segmentation in neuroscience.
+Results.
 We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples.
@@
 Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$.
-We use the log-likelihood of the autoregressive model in Keypoint-MoSeq as its prediction score and the negative energy function $f(\psi(s,a))^{\top}u(t)$ as our prediction score, respectively.
-We repeat the experiment 10 times and report the results as a box plot in Fig.
+With respect to the choice of the prediction score, for Keypoint-Moseq, we use the action log-likelihood.
+For MCD, we use the negative energy function $\phi(s,a)^{\top}u(t)$.
+For SemiSeg, we assume action variance=1 and use the Gaussian log-likelihood of actions.
+For OPAL, we use the action log-likelihood.
+We repeat the experiment 10 times and report the results as box plots in Fig.
 4A.
-Our model achieves higher AUC scores on both training and test sets (paired t-test, $p_{train}=0.0037,p_{test}=0.0046$), demonstrating a stronger ability to distinguish positive from negative samples.
-This suggests that SKIL more accurately captures time-varying pose dynamics through smoothly evolving motor primitives, while Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity.
-Beyond quantitative comparisons, we also qualitatively visualize and interpret the key skills $\phi(s,a)$ associated with example pose dynamics.
-We examine the time-varying weights $u(t)$ and compare them with the behavioral syllable segmentation produced by Keypoint-MoSeq (Fig.
+Our model achieves the highest AUC on both training and test sets (paired t-test, $p<0.05$ for every baseline), demonstrating the strongest ability to distinguish positive from negative samples.
+This suggests that MCD accurately captures time-varying pose dynamics through smoothly evolving motifs, while other models fail.
+Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity, and the autoencoders in the other two models have weaker expressive ability.
+Beyond quantitative comparisons, we also qualitatively visualize and interpret the key motifs $\phi(s,a)$ associated with example pose dynamics.
+For MCD, we examine the time-varying weights $u(t)$ of a long animal behavior video (length=250) and choose five example animal behavior clips (length=5) to check the interpretability of the result (Fig.
 4B).
-We present five example clips of observed animal behavior, each consisting of five consecutive frames.
-Each clip is characterized by a unique combination of motor skills.
-For each clip, we show the top 1-2 most dominant skills.
-For each skill, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that skill.
-The movement semantics of each skill are labeled above the visualizations.
-A more comprehensive visualization of all learned skills is included in the Appendix.
-D.
-By combining the discovered skills with real animal behavior, we assess the interpretability of each skill in the five clips (Fig.
+For our model, each clip is characterized by a unique combination of motor motifs.
+For each clip, we show the top 1-2 most dominant motifs.
+For each motif, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that motif.
+The movement semantics of each motif are labeled above the visualizations.
+A more comprehensive visualization of all learned motifs is included in the Appendix.
+F.
+For comparison, we run other models on the same behavior video, showing the latent representations by SemiSeg (Fig.
+4C) and OPAL (Fig.
+4D).
+To show clearer results for SemiSeg and OPAL, we further run KMeans (k=10) on the first 10 PCs of the latents throughout the video and show the segmentation result at the bottom of the latent representations (Fig.
+4C, D).
+We also show the behavioral syllable segmentation produced by Keypoint-MoSeq in Fig.
+4E.
+By combining the discovered motifs with real animal behavior, we assess the interpretability of each motif in the five clips (Fig.
 4B).
-First, the right-turn behavior in clip 1 is captured by the dominance of two rightward skills (skill 1 and skill 2).
-A stronger movement at the head is reflected by a higher value of skill 1.
+First, the right-turn behavior in clip 1 is captured by the dominance of two rightward motifs (motif 1 and motif 2).
+A stronger movement at the head is reflected by a higher value of motif 1.
 Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right.
-The alternate dominance of skill 3 and skill 4 aligns well with the behavior dynamics.
-Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of skill 5 and skill 6.
-Across all skills, skill 4 appears across clips 2, 3, and 4, showing its general utility.
-The transitions and mixtures of behaviors are effectively reflected in the learned skills and their temporal weights $u(t)$.
-However, when we examine the segmentation produced by Keypoint-MoSeq and the syllable dynamics (Fig.
-4C), we find inconsistencies.
-For example, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned.
+The alternate dominance of motif 3 and motif 4 aligns well with the behavior dynamics.
+Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of motif 5 and motif 6.
+Across all motifs, motif 4 appears across clips 2, 3, and 4, showing its general utility.
+The transitions and mixtures of behaviors are effectively reflected in the learned motifs and their temporal weights $u(t)$.
+However, when we examine the segmentation and latent produced by other methods (Fig.
+4C-F), we find inconsistencies.
+In Fig.
+4C, alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one.
+In Fig.
+4D, the motif weights are too dense to interpret.
+The segmentation results at the bottom of Fig.
+4 C and D do not even have repeated behavior patterns and thus could not be seen as a reasonable behavior motif representation.
+For Keypoint-Moseq, in Fig.
+4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned.
 Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements.
 For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
-Taken together, these results show that compared with similar approaches, SKIL provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor skills.
-This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor primitives.
+Taken together, these results show that compared with similar approaches, MCD provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor motifs.
+This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor motifs.
 Figure 4: A.
 AUC on training and test sets.
-B.
-We present example learned skills, behavior clips, $u(t)$, and segmentation from Keypoint-MoSeq.
-C.
-Syllables identified by Keypoint-MoSeq (adopted from [2]).
+We take an example behavior video and run our algorithm.
+We visualize the motif weights $u(t)$ and show the representative motifs in B.
+For the baseline SemiSeg, we show the latent skills and segmentation results in C.
+For the baseline OPAL, we show the latent skills and segmentation results in D.
+Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
 5 Discussion
-We introduce the SKIL framework that advances animal behavior modeling by offering a flexible, interpretable alternative to traditional discrete behavioral segmentation methods.
-It learns task-agnostic motor primitives (skill sets) that dynamically compose into complex behaviors.
-SKIL effectively recovers meaningful skills, reconstructs reward functions, and captures nuanced behavioral dynamics, providing a principled and quantitative model of how complex policies arise from fundamental motor components.
 Several limitations remain to be addressed in future work.
-First, the accuracy of inferred skills is sensitive to input data quality, as occlusions or tracking errors can degrade performance.
-Additionally, training the energy-based models can also be computationally intensive, potentially restricting real-time or large-scale applications.
-Lastly, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "skills" and specific neural circuit dynamics still requires further experimental validation.
-Beyond advancing animal behavior research, SKIL has broader implications.
+First, the accuracy of inferred motifs is sensitive to input data quality, as occlusions or tracking errors can degrade performance.
+Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "motifs" and specific neural dynamics still requires further experimental validation.
+Ethics statement
+Beyond advancing animal behavior research, MCD has broader implications.
 Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI.
@@
 A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
-Appendix A Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
+Reproducibility Statement
+We have taken several steps to ensure the reproducibility of our results.
+All source code for model training and evaluation is included in the supplementary material, allowing independent verification and replication of our experiments.
+The complete set of hyperparameter values is documented in the appendix.
+Additionally, the data preprocessing procedures and evaluation protocols are described in the main text.
+These resources provide sufficient information for reproducing the results reported in this paper.
+Appendix A LLM Usage
+In preparing this manuscript, we employed a large language model (OpenAI ChatGPT, GPT-5) as a writing assistant.
+The model was used exclusively for polishing English grammar, improving clarity, and suggesting more natural phrasing in certain sections of the text.
+All scientific content, experimental design, analyses, and interpretations were conceived, written, and verified by the authors.
+The LLM was not used to generate original research ideas, analyses, or results.
+To ensure accuracy, all model-suggested edits were carefully reviewed and, where necessary, modified by the authors.
+Appendix B Hyperparameter Setting
+We train MCD using the following hyperparameters:
+General hyperparameters.
+  • Discount factor: $\gamma=0.99$
+  • Number of epochs: $1\times 10^{6}$
+  • Batch size: 256256
+Motif representations.
+The motif representation $\phi(s,a)\in\mathbb{R}^{d}$ and $\mu(s^{\prime})\in\mathbb{R}^{d}$ were adopted with different motif dimensions $d$ depending on the task:
+  • Gridworld: $d=64$
+  • Animal navigation: $d=128$
+  • Animal free-moving: $d=64$
+Model architectures.
+  • Discrete version: $\phi$ and $\mu$ are parameterized by one-hidden-layer neural networks (hidden size =512=512).
+  • Continuous version:
+    • $\phi$: no hidden layer
+    • $\mu$: one hidden layer (hidden size =512=512)
+    • $\psi$: no hidden layer
+Learning rates.
+  • Discrete version:
+    $\phi:1\times 10^{-3},\quad\mu:1\times 10^{-3},\quad u:3\times 10^{-4},\quad w:3\times 10^{-4}.$
+  • Continuous version:
+    $\psi:5\times 10^{-4},\quad\nu:5\times 10^{-4},\quad f:3\times 10^{-4},\quad u:3\times 10^{-4}.$
+    During testing, $u$ and $f$ are further optimized on the new sequence using gradient descent with learning rate $1\times 10^{-3}$.
+We train SemiSeg and OPAL using the following hyperparameters:
+  • Discount factor: $\gamma=0.99$
+  • Number of epochs: $1\times 10^{6}$
+  • Batch size: $4\times 250$ (4 sequences, each of length=250 because this is an RNN-based inference model)
+  • Latent dimension: $d=64$
+  • Learning rate: $1\times 10^{-4}$ (tuned to get better results)
+Besides, the following loss coefficients are shared across three models for interpretability results.
+  • Temporal smoothness Gaussian random walk loss: 1010
+  • Sparsity L1-loss: $0.1$
+Appendix C Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
 By simple algebra, we obtain the quadratic potential function,
 $P(s^{\prime}|s,a)$ $\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (9)
-The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature [6, 35] and obtain the spectral decomposition of Eq.
-(7) as
+The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2014; Rahimi & Recht, 2007) and obtain the spectral decomposition of Eq.
+6 as
 $P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (10)
@@
 Note that Eq.
-(11) needs infinite $\omega$ to calculate the expectation.
+10 needs infinite $\omega$ to calculate the expectation.
 To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
@@
 $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))\approx P(s^{\prime}|s,a).$ (16)
-Hence, Eq. (7) can, in principle, yield the skill representation introduced earlier.
-Appendix B Learned Skills in the Multi-task Gridworld Dataset
-Figure 5: A. State-action maps for all 64 skills. B. Reward weight $w$ for all 64 skills.
-We visualize all original 64 skills (Fig.
+Hence, Eq. 6 can, in principle, yield the motif representation introduced earlier.
+Appendix D Multi-task Gridworld Dataset
+D.1 Dataset
+To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
+Figure 5: A. State-action maps for all 64 motifs. B. Reward weight $w$ for all 64 motifs.
+D.2 Learned Motifs
+We visualize all original 64 motifs (Fig.
 5) introduced in Sec.
-5.1.
+4.1.
 It shows some meaningful patterns as mentioned before.
-For example, skill 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid.
+For example, motif 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid.
 Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward).
-However, Task 4 (middle-middle) didn’t use this skill and used skill 39 negatively instead.
-The complex many-to-many relationship between skills and tasks informs us of the redundancy in the original skills, which inspires us to use PCA to analyze the principal components of the skill space and simplify the skill weights.
+However, Task 4 (middle-middle) didn’t use this motif and used motif 39 negatively instead.
+The complex many-to-many relationship between motifs and tasks informs us of the redundancy in the original motifs, which inspires us to use PCA to analyze the principal components of the motif space and simplify the motif weights.
 It could be seen from the comparison between Fig.
 2D and Fig.
-5A that principal components are a less redundant description of the skill space.
-Appendix C Learned Skills in the Animal Navigation Behavior Dataset
+5A that principal components are a less redundant description of the motif space.
+Appendix E Animal Navigation Behavior Dataset
+E.1 Learned Motifs
+In the original motifs of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the motif.
+Given the redundancy of the motif sets, as in Appendix.
+D, we perform PCA to analyze the principle components of the motif space and simplify the motif representation.
+To show the effect of PCA, we plot one motif (motif 0) before (Fig.
+6A right) and after PCA (Fig.
+6B right).
+Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization.
+To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in motif 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig.
+6A left and B left).
+It could be seen straightforwardly that after PCA, the motif becomes more concentrated on several $(s,a)$ pairs.
+We calculate the mean $\mu$ and variance $\sigma$ across all motifs and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right.
+We show 80 pairs before PCA and 8 pairs after PCA.
+The number of the most important pairs in each motif is called the “effective dimension.” The effective dimension is calculated across all motifs (Fig.
+6C).
+Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA.
+So the map becomes more distinct and functionally separated.
 Figure 6: A.
 Left.
-The value for each $(s,a)$ pair in skill 0 before PCA.
+The value for each $(s,a)$ pair in motif 0 before PCA.
 Right.
@@
 Left.
-The value for each $(s,a)$ pair in skill 0 after PCA.
+The value for each $(s,a)$ pair in motif 0 after PCA.
 Right.
@@
 Boxplot for the effective dimensions before and after PCA.
-In the original skills of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the skill.
-Given the redundancy of the skill sets, as in Appendix.
-B, we perform PCA to analyze the principle components of the skill space and simplify the skill representation.
-To show the effect of PCA, we plot one skill (skill 0) before (Fig.
-6A right) and after PCA (Fig.
-6B right).
-Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization.
-To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in skill 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig.
-6A left and B left).
-It could be seen straightforwardly that after PCA, the skill becomes more concentrated on several $(s,a)$ pairs.
-We calculate the mean $\mu$ and variance $\sigma$ across all skills and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right.
-We show 80 pairs before PCA and 8 pairs after PCA.
-The number of the most important pairs in each skill is called the “effective dimension.” The effective dimension is calculated across all skills (Fig.
-6C).
-Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA.
-So the map becomes more distinct and functionally separated.
-Appendix D Learned Skills in the Animal Free-moving Behavior Dataset
-Figure 7: A. Policy weight $u(t)$ for all 64 skills. B. The motion field for all 64 skills learned from the video, computed by averaging the states and actions that most strongly activate each skill.
-We show all skills learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec.
-5.3.2.
-We have completed the visualization of those skills that were previously omitted due to their perceived lack of importance.
-Due to the increased number of displayed skills, we had to renumber each skill.
-We show the present skill number above the skill motion field figure, and previous numbers (if applicable) in the parentheses following the present number.
+Appendix F Animal Free-moving Behavior Dataset
+F.1 Dataset and training
+We split the full dataset into training and test trajectories in an 8:2 ratio.
+We first learn both $f$ and $u(t)$ on the training set.
+Then, given the learned $f$, we estimate $u(t)$ on the test set.
+Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
+F.2 Learned Motifs
+We show all motifs learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec.
+4.3.
+We have completed the visualization of those motifs that were previously omitted due to their perceived lack of importance.
+Due to the increased number of displayed motifs, we had to renumber each motif.
+We show the present motif number above the motif motion field figure, and previous numbers (if applicable) in the parentheses following the present number.
+Figure 7: A. Policy weight $u(t)$ for all 64 motifs. B. The motion field for all 64 motifs learned from the video, computed by averaging the states and actions that most strongly activate each motif.

### v3

@@
 By providing a generative account of how complex animal behaviors emerge from dynamic combinations of fundamental motor motifs, our approach advances the quantitative study of natural behavior.
+behavioral neuroscience, animal behavior, compositional behavior modeling, behavior segmentation, neuroethology
 1 Introduction
 A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings.
-Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu & Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables.
-Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu & Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) switching-dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022).
+Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables.
+Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu and Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) hidden-Markov-model(HMM) based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022).
 The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
-However, existing behavior segmentation methods face four major limitations.
-First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions.
-Second, they often extract actions as abstract syllables that fail to capture how individual body parts contribute to different motions.
+However, existing behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) overlook several features of the behavior data.
+First, continuity, they model continuous behavior as combinations of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions.
+Therefore, they may fail to capture the details of delicate behavior dynamics.
+Second, compositionality, they often extract complex coordinated body movements as abstract syllables that fail to capture how individual body parts contribute to different motions.
+Therefore, they fail to reveal the connections and distinctions between syllables.
 For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns.
-Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior.
-Third, most segmentation models are either non-generative (e.g., clustering (Hsu & Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors.
-Fourth and most importantly, classification/clustering-based methods ignore temporal dependencies and fail to capture motifs reflecting true pose dynamics, while dynamics-based methods impose linear or nonlinear dynamics assumptions that may not match actual animal behavior, making such assumptions potentially harmful.
+Third, long-term dependency.
+They either ignore temporal dependency between actions or only consider a very short time window when encoding the behavior.
+Therefore, they often fail to capture the long-term and multi-scale property of syllables (See Appendix K for discussions on temporal dependency).
+Apart from these explicit features, most models are either non-generative (e.g., clustering (Hsu and Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors.
 To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework.
@@
 By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation.
-Note that Aldarondo et al.
-(2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
-We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash & Hochner, 2005).
+Note that (Aldarondo et al., 2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
+We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash and Hochner, 2005).
 Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics.
-Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2014; Ren et al., 2023; Shribak et al., 2024).
+Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2017; Ren et al., 2022; Shribak et al., 2024).
 These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts.
@@
 Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches.
-Note that there is a rich literature in robotics on learning skills through imitation learning (Paraschos et al., 2013; Lioutikov et al., 2017; Li et al., 2017; Ajay et al., 2021; Peng et al., 2022; Kuang et al., 2025).
+Further, theoretically the horizon in RL is infinite, enabling longer temporal dependency modeling.
+Note that there is a rich literature in robotics on learning skills through reinforcement learning (Appendix I).
 We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper.
@@
 They can faithfully characterize behavioral dynamics without the mismatch issues of prior segmentation methods.
-  • Our method reveals the continuous nature of animal behavior, moving beyond the discrete segmentation assumptions of existing methods.
-  • It provides a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
+  • Our method reveals the continuous, compositional and long-term dependent nature of animal behavior, providing a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
+It moves beyond the discrete segmentation assumptions of existing methods.
 2 Preliminaries
@@
 From the above definition, we can establish the following Bellman relationship:
-$Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
+$Q_{h}(s,a)$ $=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],$
+$V_{h}(s)$ $=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
 Offline Imitation Learning.
@@
 In this view, motifs are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
-Definition 1 (Motif Set).
+Definition 3.1 (Motif Set).
 Given an arbitrary transition kernel $P(\cdot|s,a)$ in an MDP, we can express it via a spectral decomposition:
@@
 The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
-Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2023; Zhang et al., 2022; Shribak et al., 2024).
+Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2022; Zhang et al., 2022; Shribak et al., 2024).
 We adopt this approach here to define motifs given the transition kernel and define rewards accordingly.
 Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
-Given the motif definition, substituting Eq. 2 and the linearized reward model into the Bellman equation (Eq. 1), we get:
-$Q(s,a)=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (3)
+Given the motif definition, substituting Eq. 2 and the linearized reward model into Bellman equation (Eq. 2), we get:
+$Q(s,a)$ $=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
+$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]$
+$=\phi(s,a)^{\top}u,\!$ (3)
 where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$.
@@
 Under this assumption, the optimal max-entropy policy $\pi(a|s)$ can be shown to follow:
-$\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
-where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$ is the entropy.
-Proposition 1.
+$\pi(a|s)$ $=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]$
+$=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
+where the entropy $H(\pi):=-\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$.
+Proposition 3.2.
 The policy in Eq.
@@
 From Def.
-1 and Prop.
-1, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state.
+3.1 and Prop.
+3.2, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state.
 Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
-Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning that characterizes the continuous behavioral dynamics) from demonstrations.
+Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning) from demonstrations.
 The learning procedure differs depending on the nature of the behavior data (i.e.
@@
 Motif discovery.
-For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) (Golub & Reinsch, 1971; Golub & Van Loan, 2013; Trefethen & Bau, 2022) or spectral decomposition representation (Ren et al., 2023; HaoChen et al., 2021) to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$.
+For discrete state-action spaces, there are two ways to learn the representation: (1) spectral methods such as singular value decomposition (SVD) (Golub and Reinsch, 1971; Golub and Van Loan, 2013; Trefethen and Bau, 2022); (2) spectral decomposition representation (Ren et al., 2022; HaoChen et al., 2021) with the objective $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$.
+Here we choose the second method and set the auxiliary distribution as $q(s^{\prime})=\rho(s^{\prime})=\int\tau^{e}(s,a)da$, i.e.
+the marginalized state distribution.
 The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
 Motif-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. 3, using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
-$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
+$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]$
+$=$ $\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
 3.2 Continuous Version
@@
 Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq.
-5 is easy to compute.
+3.1 is easy to compute.
 But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate.
-In light of these challenges, we adopt an alternative approach to learn the motif representations and policy in continuous state-action spaces.
+In light of these challenges, we adopt an alternative approach to learn the motifs and policy in continuous state-action spaces.
 Motif discovery. We model $P(s^{\prime}|s,a)$ as an energy-based model (EBM) (Shribak et al., 2024):
-$P(s^{\prime}|s,a)=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),Z(s,a)=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
+$P(s^{\prime}|s,a)$ $=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),$
+$Z(s,a)$ $=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
 where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{g}$ and $\nu:\mathcal{S}\to\mathbb{R}^{g}$ are neural-network feature maps.
@@
 2), this EBM formulation yields smooth, normalized probabilities and stable gradients, leading to more effective and generalizable motif representations.
-Proposition 2 (Connection to Motif Definition).
+Proposition 3.3 (Connection to Motif Definition).
 Given the EBM model in Eq.
-6, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
-Appendix. C contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
-To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Ma & Collins, 2018; Gutmann & Hyvärinen, 2010; 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function.
+3.2, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
+Appendix D contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
+To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Wu et al., 2018; Ma and Collins, 2018; Gutmann and Hyvärinen, 2010, 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function.
 In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq.
-6 by solving
-$\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (7)
-where $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from distribution $\rho$.
+3.2 by solving
+$\max_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{F(s,a,s^{\prime})}{F(s,a,s^{\prime})+\Sigma_{i=1}^{k}F(s,a,s^{\prime\prime}_{i}))}\right],$ (7)
+where $F(s,a,s^{\prime})=\exp(\psi(s,a)^{\top}\nu(s^{\prime}))$, $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from marginalized state distribution $\rho(s)=\int\tau^{e}(s,a)da$.
 Connection to behavioral dynamics.
-With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq.
-6.
+With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\cdot\exp\left(\|\psi(s,a)\|^{2}/2\right)\cdot\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\cdot\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq.
+3.2.
 By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier.
 Thus, Eq.
-6 offers a more general framework that extends traditional dynamics models for studying behavior.
+3.2 offers a more general framework that extends traditional dynamics models for studying behavior.
 Moreover, in practice, using Eq.
-6 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data.
+3.2 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data.
 Thus, while Eq.
 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq.
-6 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
+3.2 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
 Motif-based policy learning.
 After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq.
-7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix.
-C).
+7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix D).
 However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly.
@@
 As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq.
-5 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space.
+3.1 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space.
 This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$.
 Thus, it’s reasonable to apply NCE here again,
-$\min_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{\exp(f(\psi(s,a))^{\top}u)}{\exp(f(\psi(s,a))^{\top}u)+\Sigma_{i=1}^{k}\exp(f(\psi(s,a^{\prime}_{i}))^{\top}u)}\right],$ (8)
-where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
-3.3 Understanding Animal Behavior via Motif and Policy Learning
-In this section, we discuss how motor motifs $\phi$ and motif weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u$ to vary across tasks $t$ or time points $t$.
-As the motif coefficients, $u(t)$ would dynamically modulate the influence of each motif on the final policy.
-We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
-The concrete results will be later shown in Sec.
-4.2 and Sec.
-4.3 respectively.
-In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task or time point.
-Scenario (1): Consider a mouse navigating a maze (Rosenberg et al., 2021), where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay).
-We assume the animal switches between $T$ strategies, each associated with a unique reward, that guide its navigation, with the timing of each reward condition known from Ke et al.
-(2025).
+$\max_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{G(s,a)}{G(s,a)+\Sigma_{i=1}^{k}G(s,a^{\prime}_{i})}\right],$ (8)
+where $G(s,a)=\exp(f(\psi(s,a))^{\top}u)$, $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples. Identifiability discussions are provided in Appendix J.
+3.3 Understanding Animal Behavior via Motifs
+In this section, we discuss how motifs $\phi$ and weights $u$ can be used to describe animal behavior trajectories, by allowing $u$ to vary across tasks or time points and dynamically modulate each motif.
+We consider two scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
+In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task $t$ or time point $t$.
+Scenario (1): Consider a mouse navigating a maze, where trajectories are discretized into a finite state and action space.
+We assume the animal switches between $T$ strategies, each associated with a unique reward function.
 We first learn shared motifs $\phi(s,a)$ using Eq.
 2, then fit task-specific policies $\pi(a|s,t)$ using Eq.
-5, where each task $t$ corresponds to one of $T$ strategies.
-This yields $T$ sets of weights $u(t)$, one per task, while sharing a common motif set across tasks.
-Scenario (2): A representative case is a freely behaving mouse (Wiltschko et al., 2015; Weinreb et al., 2024), where the state is defined by pose keypoints and the action by state change–both continuous.
-We first learn $\psi(s,a)$ from pose trajectories using Eq.
-7.
-Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq.
-8, yielding the motor motif $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$.
-Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
-Beyond capturing smoothly time-varying motif compositions, our framework allows multiple motifs to be active simultaneously.
-Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl(\phi(s,a)^{\top}u(t)\bigr)$, each action is a generalized linear combination of basis motifs weighted by $u(t)$, enabling overlapping and composable behaviors.
-For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—recruiting different motifs concurrently.
-Unlike discrete switching-state models, which assign one behavior per state, our continuous motif-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
-In Sec.
-4, we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
+3.1, where each task $t$ is one of the $T$ strategies.
+This yields $T$ sets of weights $u(t)$, one per task, which share a common motif set $\phi(s,a)$ across tasks.
+Scenario (2): Consider a mouse freely moving in the cage, where the states are the continuous coordinates of the body parts, and actions are the continuous velocities.
+We learn $\psi(s,a)$ from pose trajectories using Eq.
+7, and learn $u(t)$ and $f$ using Eq.
+8.
+Finally, we get motor motifs in $\phi(s,a)=f(\psi(s,a))$, and how they can construct time-varying policy $\pi(a|s,t)$ through weights $u(t)$.
+The policy is finally represented by a generalized linear combination of basis motifs, producing two advantages: (1) Continuity.
+Unlike models with abrupt discrete switches, our approach features continuous-time formulation with smoothly evolving motifs, and can capture subtle behavioral changes more faithfully; (2) Compositionality.
+Unlike models assigning one discrete label to one timestep, our approach provides a compositional view and allow two motifs to be active simultaneously.
+For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side.
+(3) Long-term dependency.
+Unlike models only considering short time windows, our approach has multi-scale temporal dependency and infinite horizon in theory (Appendix K).
+Taken together, MCD offers a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a continuous, compositional and multi-scale time-varying description of animal trajectories.
+In Experiment (Sec.
+4), we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
 3.4 Reward Recovery
@@
 This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals.
-Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008), where both the policy and underlying reward are inferred from demonstrations.
-In the context of animal behavior (Ke et al., 2025; Zhu et al., 2024; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes.
-Since $V(s,t)$ can only be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is discrete and finite.
+Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008).
+In the context of animal behavior (Ke et al., 2025; Zhu et al., 2023; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal decision-making processes.
+Since $V(s,t)$ can only be easily computed in closed form in discrete and finite settings, we validate our method by visualizing the inferred rewards in the first two experiments.
 4 Experiments
-4.1 Application to simulated data in a multi-task gridworld
-The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig.
-2A).
-In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$.
-Fig.
-2C (left) shows the ground truth of reward functions for all nine tasks.
-In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$.
-See Appendix.
-D for more details for data generation.
-As described in Sec.
-3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$.
-Because the computational complexity only scales linearly with the number of motifs, we assume a total of 64 motifs and learn the 64 $\phi$ vectors to cover the motif space as much as possible.
-Visualizations of these motifs are shown in the Appendix.
-D.
-Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec.
-3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
-This recovered reward (Fig.
-2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
-To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig.
-2B).
-Therefore, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig.
-2D, treating them as basis vectors spanning the motif space.
-The PC motifs exhibit interpretable patterns.
-For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values.
-This motif corresponds to moving away from the middle-right grid toward the bottom-left.
-Examining the $w$ matrix, we see that motif 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task.
-In contrast, motif 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal.
-Similar interpretations can be made for other motifs and tasks.
-In this gridworld experiment, we successfully recover reward functions from behavior trajectories, and, importantly, the learned $\phi$ and $w$ are effectively deployed in different task settings.
+4.1 Application to simulated multi-task gridworld
 Figure 2: A.
@@
 In the PC motif plot, red indicates positive feature values, blue indicates negative.
+The gridworld consists of a $3\times 3$ lattice, and each of the nine states allows four possible actions: Up, Down, Left, and Right (Fig.
+2A).
+In task $t\in\{0,...,8\}$, a fixed reward is assigned to the $(s,a)$ pairs that move toward state $t$.
+Fig.
+2C (left) shows the ground truth of reward functions for all nine tasks.
+We collect a dataset where in each episode, the agent starts from a random state and navigates to task-specific state $t$.
+See Appendix E for more details for data generation.
+As described in Sec.
+3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$.
+Because the computational complexity only scales linearly with the number of motifs, to cover the motif space as much as possible, we select a large number for motif dimension $d=64$.
+(See Appendix L.1 for discussions of motif dimensions.) Visualizations of these motifs are shown in the Appendix E.
+Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec.
+3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
+This recovered reward (Fig.
+2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
+To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig.
+2B).
+These components are basis vectors spanning the motif space.
+To check if they exhibit interpretable patterns, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig.
+2D.
+For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values.
+This motif corresponds to moving away from the middle-right grid toward the bottom-left.
+Accordingly, we observe at Column 0 in $w$ a high value at Row 6 and a low value at Row 5, suggesting that Task 6 requires this motif to construct its reward function.
+Similar interpretations can be made for other motifs and tasks.
+Therefore, the learned motifs are effectively deployed to recover reward functions.
 4.2 Application to animal navigation behavior
 Dataset and model setup.
-We next evaluate our method on a real animal behavior dataset from Rosenberg et al.
-(2021).
-In this experiment, a thirsty mouse is trained to navigate in a binary-tree maze (Fig.
-3A), starting each trial from the central home cage and attempting to reach a water port at one leaf node.
+We next evaluate our method on a real animal behavior dataset from (Rosenberg et al., 2021).
+In this experiment, a thirsty mouse is trained to start from the home cage and navigate to the water port in a binary-tree maze (Fig.
+3A).
 The state space is defined by the mouse’s location on the tree.
 The actions include moving to its left parent, right parent, left child, and right child.
-Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter.
-This complex behavior cannot be captured by a single reward function.
-Studying this dataset allows us to discover motifs shared across multiple reward functions and policies.
-This in turn tests whether complex navigation behavior, under multiple competing motivations, can be distilled into a small set of reusable motifs that provide interpretable insight into animal decision-making.
-We first apply the segmentation algorithm from Ke et al.
-(2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration.
-The algorithm could also infer reward maps for these tasks, which are shown in Fig.
-3A (top row).
+Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter, which cannot be simply captured by a single reward function.
+Studying this dataset allows us to discover reusable and interpretable motifs shared across multiple competing motivations.
+We first apply the segmentation algorithm from (Ke et al., 2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration, and infer their respective reward functions (Fig.
+3A, top row).
 While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior.
-Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting.
-To uncover a task-agnostic set of motifs capable of constructing all three reward functions, we apply our model as described in Sec.
-3.1.
-Again, to cover the motif space as much as possible, we set the number of motifs $d=128$.
+Then we apply our model to identify motifs shared across these three reward functions (discrete version, Sec.
+3.1).
+Again, we choose the motif dimension $d=127$ to completely cover the motif space.
+See Appendix L.1 for more details on the selection of $d$.
 Figure 3: A.
@@
 Results.
-To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
+To assess model performance, we estimate the task-specific weights $w(t)=u(t)-\gamma\int V(s^{\prime},t)\mu(s^{\prime})q(s^{\prime})ds^{\prime}$ (Eq.
+3) and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$.
 The recovered reward functions align closely with the ground truth (Fig.
@@
 In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it.
-During home-seeking, a distinct peak appears at the home cage.
-When exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
+In home-seeking, a distinct peak appears at the home cage.
+In exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
 We further visualize the learned motifs by applying PCA to obtain PC motifs.
-See the Appendix.
-E for raw features of motifs.
+See Appendix F for all raw motifs.
 Fig.
-3B displays two top-contributing motifs for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig.
+3B displays two top-contributing motifs for the water and home task, selected based on the peak of linear weights $w$ (Fig.
 3C).
-The water-related motifs promote movement toward the water port, while the home-related motifs guide navigation to the cage.
-Notably, the motifs important for one task have minimal or negative contributions to the other (Fig.
-3C), indicating clear functional specialization.
-Fig.
-3C also shows non-task-specific motifs with similar weights in both water- and home-seeking tasks.
-These results show that our model not only recovers multiple reward functions from real behavior but also learns interpretable motifs whose contributions to each task are distinct and behaviorally meaningful.
-Unlike previous reward discovery on this dataset (Ashwood et al., 2022; Ke et al., 2025), which cannot identify such motifs, our approach reveals how reward maps can be decomposed into smaller, reusable action components.
-This decomposition offers a mechanistic view of how local decision processes combine to produce strategies such as water-seeking or home-seeking.
+The water motifs promote movement towards the water port, while the home motifs promote movement towards the cage.
+Moreover, most motifs are dominated by one task rather than shared across them (Fig.
+3C), reflecting competitive motivations.
+Unlike previous reward discovery methods on this dataset that lack motifs  (Ashwood et al., 2022; Ke et al., 2025), these results show that our model can learn one single set of interpretable motifs to recover multiple reward functions, revealing how local decision processes combine to produce complex strategies.
 4.3 Application to animal free-moving behavior
+Figure 4: A.
+Human annotated label prediction accuracy and AUC on test sets.
+AUC experiments are repeated for 10 sequences with 100 timesteps for each sequence.
+The negative samples are sampled 1000 times for each sequence.
+We take an example behavior video and run our algorithm.
+We visualize the motif weights $u(t)$ and show the representative motifs in B.
+For the baseline SemiSeg, we show the latent skills and segmentation results in C.
+For the baseline OPAL, we show the latent skills and segmentation results in D.
+Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
 Dataset and model setup.
-We finally apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024) to extract motor motifs and analyze fine-grained pose dynamics.
-This dataset contains the keypoint coordinates of eight mouse body parts, including the head, the nose, both ears, and four spine nodes.
-Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part.
-The action at each timestep is defined as the velocity of state, $a_{t}=(s_{t+1}-s_{t})/\delta t$.
-We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset).
-We formulate the data using a continuous MDP.
-Studying this dataset allows us to ask whether free-moving behaviors, which often appear as mixtures of grooming, locomotion, and postural adjustments, can be represented as combinations of a compact set of task-agnostic motor motifs.
-As outlined in Sec.
-3.2, we use NCE to learn the motif representation $\phi(s,a)$ and the time-dependent weights $u(t)$.
-To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$.
-To perform more stable optimization, we optimize $\phi(s,a)$ and $u(t)$ using coordinate descent, updating them alternately.
-For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses.
-Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned motif representation $\phi$ and the temporal weights $u(t)$.
-With respect to the choice of the number of motifs, we find in practice that the performance grows more slowly once past $d=64$, so we choose this as an optimal number.
-See Appendix.
-F for more training details.
-We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), as a representative for switching-dynamics-based segmentation methods; and (2) SemiSeg (Whiteway et al., 2021), as a representative for clustering-based behavior segmentation methods.
-Note that Keypoint-MoSeq is regarded as a SOTA approach, because it extends the autoregressive hidden Markov model (AR-HMM) and MoSeq (Wiltschko et al., 2015), and has been shown to outperform B-SOiD (Hsu & Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014).
-We also include OPAL (Ajay et al., 2021), a representative autoencoder-based motif learning algorithm from robotics, as a baseline to highlight the advantages of our method and why robotics approaches are ill-suited for behavioral segmentation in neuroscience.
+To verify MCD’s generalizability, we apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024).
+The state corresponds to the $x,y$ coordinate of eight body parts, including the head, nose, both ears, and four spine nodes.
+The action is defined as the body part velocity, $a_{t}=(s_{t+1}-s_{t})/\delta t$.
+We set the time interval to $\delta t=1$ ($1/30s$ in the real world).
+Studying this dataset allows us to ask whether free-moving behaviors, which consist of mixtures of grooming, locomotion, and postural adjustments, can be combined through a compact set of motifs.
+For this continuous dataset, we use NCE to learn the motifs $\phi(s,a)$ (Sec.
+3.2) and the time-dependent weights $u(t)$ (Sec.
+3.3).
+We place a Gaussian random walk prior on $u(t)$ to ensure temporal smoothness: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$.
+We optimize $f$ and $u(t)$ using coordinate descent to improve stability.
+Since here our focus is on understanding the learned motifs and animal dynamics, we do not perform IRL but focus on interpreting $\phi$ and $u(t)$.
+We find in practice that the performance grows slowly once the motif dimension exceeds $d=64$ (Appendix L.1), so we choose this in practice.
+See Appendix G for more training details.
+We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), an HMM-based method; and (2) SemiSeg (Whiteway et al., 2021), a clustering-based method.
+Note that we only select one HMM-based method, since Keypoint-MoSeq methodologically extends the autoregressive hidden Markov model (AR-HMM, as in MoSeq (Wiltschko et al., 2015)), and is the state-of-the-art approach, outperforming a series of work (B-SOiD (Hsu and Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014)).
+We also include OPAL (Ajay et al., 2021), an autoencoder-based motif learning algorithm from robotics, as a baseline to show our advantages, and that robotics approaches are ill-suited for behavioral segmentation in neuroscience.
 Results.
-We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples.
-AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score.
+To evaluate whether the inferred latents $u(t)$ are meaningful, we fit a two-layer neural network to decode the human-annotated labels.
+MCD achieves the highest test accuracy (Fig.
+4A top), suggesting motifs recovered by MCD align better with actual animal behavior dynamics.
+More details are provided in Appendix H.
+To evaluate the model’s ability to discriminate real trajectories from mismatch trajectories, we calculate the area under the receiver operating characteristic curve (AUC), which allows direct comparison between our unnormalized energy function and other models’ normalized likelihood.
 Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$.
-With respect to the choice of the prediction score, for Keypoint-Moseq, we use the action log-likelihood.
-For MCD, we use the negative energy function $\phi(s,a)^{\top}u(t)$.
-For SemiSeg, we assume action variance=1 and use the Gaussian log-likelihood of actions.
-For OPAL, we use the action log-likelihood.
-We repeat the experiment 10 times and report the results as box plots in Fig.
-4A.
-Our model achieves the highest AUC on both training and test sets (paired t-test, $p<0.05$ for every baseline), demonstrating the strongest ability to distinguish positive from negative samples.
-This suggests that MCD accurately captures time-varying pose dynamics through smoothly evolving motifs, while other models fail.
-Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity, and the autoencoders in the other two models have weaker expressive ability.
-Beyond quantitative comparisons, we also qualitatively visualize and interpret the key motifs $\phi(s,a)$ associated with example pose dynamics.
-For MCD, we examine the time-varying weights $u(t)$ of a long animal behavior video (length=250) and choose five example animal behavior clips (length=5) to check the interpretability of the result (Fig.
+For the choice of the prediction score, we use (1) action log-likelihood for Keypoint-MoSeq; (2) negative energy function $\phi(s,a)^{\top}u(t)$ for MCD; (3) Gaussian log-likelihood ($\sigma=1$) for SemiSeg; and (4) action log-likelihood for OPAL.
+Our model achieves the highest AUC (paired t-test, $p<0.05$ for every baseline, Fig.
+4A bottom), demonstrating the strongest ability to distinguish real trajectories and accurately capture continuous dynamics through compact motifs.
+Beyond quantitative comparisons, we also qualitatively visualize and interpret $\phi(s,a)$ associated with example pose dynamics.
+For MCD, we examine the time-varying $u(t)$ of a long animal behavior video (length=250), from which we extract five clips (length=5), each of which corresponds to a unique mixture of motifs (Fig.
 4B).
-For our model, each clip is characterized by a unique combination of motor motifs.
-For each clip, we show the top 1-2 most dominant motifs.
-For each motif, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that motif.
-The movement semantics of each motif are labeled above the visualizations.
-A more comprehensive visualization of all learned motifs is included in the Appendix.
-F.
-For comparison, we run other models on the same behavior video, showing the latent representations by SemiSeg (Fig.
+For each clip, we show the top 1-2 most dominant motifs, and display them using average pose and actions, computed by averaging the states and actions that most strongly activate that motif.
+Each motif is labeled with its meaning.
+A more comprehensive visualization is included in Appendix G.
+On the same video, we show the latents inferred by SemiSeg (Fig.
 4C) and OPAL (Fig.
 4D).
-To show clearer results for SemiSeg and OPAL, we further run KMeans (k=10) on the first 10 PCs of the latents throughout the video and show the segmentation result at the bottom of the latent representations (Fig.
-4C, D).
-We also show the behavioral syllable segmentation produced by Keypoint-MoSeq in Fig.
+For clarity, we further run KMeans (k=10) on the first 10 PCs of the latents and show the clustering result at the bottom.
+The segmentation produced by Keypoint-MoSeq is shown in Fig.
 4E.
@@
 The transitions and mixtures of behaviors are effectively reflected in the learned motifs and their temporal weights $u(t)$.
-However, when we examine the segmentation and latent produced by other methods (Fig.
-4C-F), we find inconsistencies.
-In Fig.
-4C, alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one.
-In Fig.
-4D, the motif weights are too dense to interpret.
-The segmentation results at the bottom of Fig.
-4 C and D do not even have repeated behavior patterns and thus could not be seen as a reasonable behavior motif representation.
+However, other models show inconsistencies.
+In SemiSeg (Fig.
+4C), alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one.
+OPAL (Fig.
+4D) lacks dominant and specific motifs for interpretation.
+Moreover, these two models do not even have segmentation label that appears again, suggesting their latents can not decompose behaviors to a simple subset.
 For Keypoint-MoSeq, in Fig.
@@
 For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
-Taken together, these results show that compared with similar approaches, MCD provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor motifs.
-This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor motifs.
-Figure 4: A.
-AUC on training and test sets.
-We take an example behavior video and run our algorithm.
-We visualize the motif weights $u(t)$ and show the representative motifs in B.
-For the baseline SemiSeg, we show the latent skills and segmentation results in C.
-For the baseline OPAL, we show the latent skills and segmentation results in D.
-Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
+Taken together, these results show that compared with baselines, MCD could accurately capture complex behaviors through a compact, task-agnostic set of motor motifs, offering a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motifs.
+We provide ablation studies in Appendix C to show that MCD is robust to motif dimension $d$, policy noise distribution, dataset partition, and Gaussian random walk prior.
+We also evaluate the generative quality of MCD by performing several rollouts in Appendix 16.
 5 Discussion
+We talk about the relations of our work to neuroscience and ethology in Appendix M.
 Several limitations remain to be addressed in future work.
-First, the accuracy of inferred motifs is sensitive to input data quality, as occlusions or tracking errors can degrade performance.
-Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "motifs" and specific neural dynamics still requires further experimental validation.
-Ethics statement
+First, the accuracy of inferred motifs is sensitive to input data quality, and tracking errors can degrade performance.
+Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned ”motifs” and specific neural dynamics still requires further experimental validation.
+Impact statement
 Beyond advancing animal behavior research, MCD has broader implications.
@@
 A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
-Reproducibility Statement
+Appendix A Reproducibility Statement
 We have taken several steps to ensure the reproducibility of our results.
@@
 These resources provide sufficient information for reproducing the results reported in this paper.
-Appendix A LLM Usage
+Appendix B LLM Usage
 In preparing this manuscript, we employed a large language model (OpenAI ChatGPT, GPT-5) as a writing assistant.
@@
 To ensure accuracy, all model-suggested edits were carefully reviewed and, where necessary, modified by the authors.
-Appendix B Hyperparameter Setting
+Appendix C Hyperparameter Setting
 We train MCD using the following hyperparameters:
@@
   • Gridworld: $d=64$
-  • Animal navigation: $d=128$
+  • Animal navigation: $d=127$
   • Animal free-moving: $d=64$
@@
   • Continuous version:
-    • $\phi$: no hidden layer
-    • $\mu$: one hidden layer (hidden size =512=512)
-    • $\psi$: no hidden layer
+    • $f$: no hidden layer
+    • $\nu$: one hidden layer (hidden size =512=512), with one normalization layer to ensure the L2-Norm of the output is 1.
+    • $\psi$: no hidden layer, with one normalization layer to ensure the L2-Norm of the output is 1.
+  • In both cases, $u$ and $w$ are just two matrices, with each column corresponding to $u(t)$ or $w(t)$ at a specific task or timepoint.
+  • Activation function: For all networks, if it has at least one hidden layer, then all of its activation functions are Exponential Linear Unit (ELU).
 Learning rates.
   • Discrete version:
-    $\phi:1\times 10^{-3},\quad\mu:1\times 10^{-3},\quad u:3\times 10^{-4},\quad w:3\times 10^{-4}.$
+    $\phi:1\times 10^{-3},\mu:1\times 10^{-3},u:3\times 10^{-4},w:3\times 10^{-4}.$
   • Continuous version:
-    $\psi:5\times 10^{-4},\quad\nu:5\times 10^{-4},\quad f:3\times 10^{-4},\quad u:3\times 10^{-4}.$
+    $\psi:5\times 10^{-4},\nu:5\times 10^{-4},f:3\times 10^{-4},u:3\times 10^{-4}.$
     During testing, $u$ and $f$ are further optimized on the new sequence using gradient descent with learning rate $1\times 10^{-3}$.
@@
   • Sparsity L1-loss: $0.1$
-Appendix C Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
-By simple algebra, we obtain the quadratic potential function,
-$P(s^{\prime}|s,a)$ $\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (9)
-The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2014; Rahimi & Recht, 2007) and obtain the spectral decomposition of Eq.
-6 as
+Appendix D Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
+In this part, we show in detail the connection between the EBM formulation (Eq. 3.2) and the low-rank spectral decomposition formulation (Eq. 2) of the transition kernel $P(s^{\prime}|s,a)$.
+From Eq. 3.2, by simple algebra, we obtain the quadratic potential function,
+$P(s^{\prime}|s,a)\propto$ $q(s^{\prime})\cdot\exp\left(\|\psi(s,a)\|^{2}/2\right)\cdot\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$
+$\cdot\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right).$ (9)
+The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2017; Rahimi and Recht, 2007) and obtain the spectral decomposition of Eq.
+3.2 as
 $P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (10)
 where $\omega\sim\mathcal{N}(0,I)$ is the frequency in the Fourier domain, and
-$\phi_{\omega}(s,a)$ $=\exp\left(-i\omega^{\top}\psi(s,a)\right)\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (11)
-$\mu_{\omega}(s^{\prime})$ $=\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (12)
+$\phi_{\omega}(s,a)=$ $\exp\left(-i\omega^{\top}\psi(s,a)\right)$
+$\cdot\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (11)
+$\mu_{\omega}(s^{\prime})=$ $\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\cdot\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (12)
 Note that Eq.
@@
 Then it’s straightforward to see that,
-$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))\approx P(s^{\prime}|s,a).$ (16)
-Hence, Eq. 6 can, in principle, yield the motif representation introduced earlier.
-Appendix D Multi-task Gridworld Dataset
-D.1 Dataset
-To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
+$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ $=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))$
+$\approx P(s^{\prime}|s,a).$ (16)
+Hence, Eq. 3.2 can, in principle, yield the motif representation introduced earlier.
+Appendix E Multi-task Gridworld Dataset
+E.1 Dataset
+To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)={\exp(Q(s,a,t))}/{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
 Figure 5: A. State-action maps for all 64 motifs. B. Reward weight $w$ for all 64 motifs.
-D.2 Learned Motifs
+E.2 Learned Motifs
 We visualize all original 64 motifs (Fig.
@@
 5A that principal components are a less redundant description of the motif space.
-Appendix E Animal Navigation Behavior Dataset
-E.1 Learned Motifs
+Appendix F Animal Navigation Behavior Dataset
+F.1 Learned Motifs
 In the original motifs of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the motif.
-Given the redundancy of the motif sets, as in Appendix.
-D, we perform PCA to analyze the principle components of the motif space and simplify the motif representation.
+Given the redundancy of the motif sets, as in Appendix E, we perform PCA to analyze the principle components of the motif space and simplify the motif representation.
 To show the effect of PCA, we plot one motif (motif 0) before (Fig.
@@
 Boxplot for the effective dimensions before and after PCA.
-Appendix F Animal Free-moving Behavior Dataset
-F.1 Dataset and training
+Appendix G Animal Free-moving Behavior Dataset
+G.1 Dataset and training
 We split the full dataset into training and test trajectories in an 8:2 ratio.
@@
 Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
-F.2 Learned Motifs
+G.2 Learned Motifs
 We show all motifs learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec.
@@
 Figure 7: A. Policy weight $u(t)$ for all 64 motifs. B. The motion field for all 64 motifs learned from the video, computed by averaging the states and actions that most strongly activate each motif.
+Appendix H Evaluation details on a Human-Annotated Dataset.
+We verify our method on a supervised behavior label dataset.
+Due to the substantial cost and labor demands of manual annotation, large-scale, high-quality animal behavior datasets with human labels remain limited.
+Commercial annotation tools (e.g., HomeCageScan) are also expensive, further constraining availability.
+To support quantitative evaluation, we therefore included a human-annotated subset from the freely moving animal behavior dataset we mentioned last section.
+Specifically, we selected 200 short video clips (10 recording sessions; in each session we randomly sampled 20 clips, with 200 frames each).
+Each clip was manually annotated with behavior segments for each frame, resulting in 200 per-frame labels across six categories: Walking, Sniffing/Grooming, Turning left, Turning right, Rearing, and Resting.
+Manual annotation required approximately 40 hours of expert effort.
+The dataset is split into a training set and a test set with a ratio $8:2$.
+As before, we evaluated four models on this dataset: MCD, Keypoint-MoSeq, SemiSeg, and OPAL.
+Because the latent variables or syllables inferred by these models do not necessarily align with human-defined labels, we trained an additional decoder to map model-specific latent representations to the ground-truth categories.
+For each model, the decoder input was the motif weights $u(t)$ (MCD), the inferred one-hot syllable (Keypoint-MoSeq), or the latent embedding $z$ (SemiSeg and OPAL).
+The decoder was implemented as a two-layer neural network with ReLU activation, optimized using the Adam optimizer and cross-entropy loss (learning rate =0.001=0.001, hidden size=100=100), following standard settings in the scientific computation package scikit-learn.
+The resulting classification accuracies on the held-out test set (Fig.
+4A) demonstrate that MCD substantially outperforms all baselines, achieving near-perfect accuracy.
+This indicates that the motifs recovered by MCD align better with actual animal behavior dynamics.
+Appendix I Connections to Relevant Algorithms in RL
+Motor primitives learning.
+Animal behavior research has focused on identifying motor primitives directly from large-scale recordings.
+Supervised classification (Marks et al., 2022; Segalin et al., 2021), clustering-based analysis (Hsu and Yttri, 2021; Berman et al., 2014), and HMM-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Whiteway et al., 2021) label behavior into discrete syllables and analyze neural correlates of transitions.
+While these approaches reveal stereotyped motifs such as walking or grooming, they typically treat primitives as rigid and exclusive, overlooking the continuous, compositional structure of behavior and its adaptation across contexts.
+In robotics, movement primitives have long been studied (Paraschos et al., 2013; Saveriano et al., 2023; Lioutikov et al., 2017).
+ProMP (Paraschos et al., 2013) extracts primitives from demonstrations but requires labeled skills/motifs, while ProbS (Lioutikov et al., 2017) jointly infers behavioral segmentation and a primitive library via Expectation-Minimization (EM), enabling unsupervised skill discovery.
+These approaches resemble dynamics-based behavior segmentation in animal behavior research.
+All methods mentioned here lack an RL perspective that links primitives to policies and reward-driven behavior.
+Our framework instead leverages offline RL-based imitation learning to uncover motif-based policies that capture both modular primitives and their sequential composition, and provides a principled way to explain why behaviors occur, not just how they unfold, and how reusable motifs/skills contribute to them in a generative decision-making process.
+Offline Imitation learning.
+Offline imitation learning presents the problem of learning a policy from fixed demonstrations when access to environments is impossible.
+Simple behavior cloning can be performed offline, but fails to generalize well in some cases because it does not consider dynamics or environment structure limits.
+ValueDICE (Kostrikov et al., 2019) considers the dynamics of the training data and learns a policy that minimizes the KL-divergence between the state-action occupancies generated by the policy and of the original dataset.
+But the adversarial optimization of policy and Q-functions introduces instability in training.
+IQL (Garg et al., 2021) avoids adversarial training by directly parameterizing the policy in terms of the $Q$-function, $\pi(a|s)=\exp(Q(s,a))/Z$.
+However, all of these algorithms learn policies as unstructured functions.
+As the animal behavior structure is highly modular and stereotyped, it is more appropriate to employ a hierarchical motif-based policy to model the data.
+Motif/Skill discovery.
+Unsupervised skill discovery (DIAYN (Eysenbach et al., 2019), BeCL (Yang et al., 2023), DADS (Sharma et al., 2020), InfoGAIL (Li et al., 2017), Directed-InfoGAIL (Sharma et al., 2018), SkillBlender (Kuang et al., 2025), ASE (Peng et al., 2022)) to find a high-level abstraction of actions has been an effective strategy for online RL and imitation learning.
+However, they are limited to online settings when the skills can only be refined through interacting with the environment, which restricts their application to large-scale offline datasets.
+Recent offline skill discovery methods include OPAL (Ajay et al., 2021), SPiRL (Pertsch et al., 2021a), and SkiLD (Pertsch et al., 2021b) which use an autoencoder to encode trajectories into latent skills $z$; and PARROT (Singh et al., 2021) which uses a flow-based model to learn a behavior prior.
+In these models, the latent $z$ is later used to generate the policy $\pi(a|s,z)$ in a non-linear way.
+Our paper, instead, employs a generalized-linear structure of the policy $\pi(a|s,z)\propto\exp(\phi(s,a)^{\top}u(z=t))$ which provides better interpretability than the policy network.
+This interpretability is essential for addressing downstream neuroscience questions.
+Linear structure of environment/policy.
+Another line of work (HILP (Park et al., 2024), FB (Touati and Ollivier, 2021), USFA (Borsa et al., 2019), RaMP (Chen et al., 2023)) based on successor features (SF) also uses a generalized linear structure to model the motif-based policy as $\pi(a|s,z)\propto\exp(F(s,a,z)^{\top}z)$ where $F(s,a,z)$ is SF under a certain motif $z$.
+The main concern is that they cannot separate motif representation $z$ from state-action representation in SF, while our work can (Eq.
+3).
+Therefore, their motifs depend on the specific task or timepoint, while our motifs/skills are general representations shared across tasks and timepoints.
+The idea that task-agnostic motifs are combined adaptively to form new policies aligns better with the need of interpretability and scientific discovery: we would like to look for neural signals responsible for relatively fixed time-agnostic behavior patterns, to help us better understand the animal behavior.
+TRAIL (Yang et al., 2022) adopts a linear decomposable environment as MCD when learning the latent skills.
+However, the latent-conditioned policy $\pi(a|s,z)$ is not linear, but parametrized by a neural network.
+SkillBlender (Kuang et al., 2025) adopts a linear decomposable policy, and uses linear combinations of lower-level controller outputs as the final actions $a_{t}=\sum_{i}k_{i}a^{i}_{t}$, while we use the linear combinations of different motifs to generate the final state-value functions $Q_{t}(s,a)=\sum_{i}\phi_{i}(s,a)u^{i}_{t}$ rather than direct actions.
+We believe a naive mixture of the low-level controller is less biologically realistic in modeling animal behavior than mixtures of state-value functions.
+The latter can find supporting evidence in neuroscientific literature (Makino, 2023).
+Besides, their transition kernel $P(s^{\prime}|s,a)$ is not linear.
+Compared to them, our work assumes generalized linear structures for both the policy and the environment, providing better interpretability.
+This shared motif is more fundamental in revealing the basic structure of the animal’s intentions.
+Our contributions.
+Most existing skill-learning methods do not apply to the behavior understanding scenario considered here.
+We study offline data without supervision or task annotations—no explicit goals, labels, or trajectory segmentation—only long, unstructured behavior sequences.
+The challenge is threefold: (1) to discover the basic motor skills, (2) to determine how these skills compose spatially and temporally within long trajectories in an offline setting, and (3) to ensure interpretability for scientific discovery.
+These are nontrivial problems beyond the reach of existing skill- or motor-primitive approaches.
+To our knowledge, this is the first work to introduce offline RL-based imitation learning for behavior segmentation, yielding interpretable skill representations and their compositions.
+Appendix J Identifiability Discussion
+In this section we clarify the notion of identifiability and how it applies to our motif-based continuous dynamics (MCD) model.
+A parameterization is said to be identifiable if different parameter values lead to different model predictions.
+Conversely, if multiple distinct parameterizations yield identical likelihoods or value functions, the model is only identifiable up to an equivalence class (e.g., scaling, rotation, or permutation).
+Our model is identifiable at the subspace level.
+In MCD, both the transition operator and the action-value function admit the spectral form
+$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime}),\qquad Q(s,a)=\phi(s,a)^{\top}u.$
+Although individual components of $\phi$, $\mu$, and $u$ may be rescaled by an invertible linear transformation without altering $P$ or $Q$, the subspace spanned by the motif functions is uniquely determined (Ren et al., 2022).
+This is analogous to PCA, where the principal subspace is identifiable even if the basis vectors within that subspace are not uniquely determined.
+In practice, they converge well.
+Moreover, in the continuous version, the identifiability of the new representations $\psi,\nu$ up to rescaling has been guaranteed by the normalization operation before the output (Appendix C), which also ensures the trainability of the contrastive learning algorithm.
+Comparison to HMMs and latent variable models.
+Classical behavioral segmentation methods (e.g., Keypoint-Moseq as an HMM variant, SemiSeg as an autoencoder variant) are generally not identifiable without strong additional assumptions.
+Different transition matrices, emission matrices, or latent dynamics parameters can produce indistinguishable observation distributions.
+These models typically suffer from label-swapping and more severe transform equivalences in the latent space, meaning that even the latent subspace is not uniquely determined.
+In contrast, our method fixes the ambient representation through the observed state–action pairs and learns a function class whose span is uniquely tied to the transition operator and value function.
+Thus, while our motif basis is identifiable only up to linear transforms (as is standard for spectral decompositions), the underlying subspace and its behavioral interpretation are stable, reproducible and identifiable.
+Summary.
+MCD inherits the identifiability properties of spectral RL methods: the motif basis is identifiable up to an invertible transformation, and the motif subspace—the structure that determines $P$, $Q$, and $\pi$—is uniquely recoverable.
+This stands in contrast to latent-variable models such as HMMs or SLDS, which in general do not admit identifiable latent subspaces without strong, often unrealistic, constraints.
+Appendix K Long-term Dependency Discussion
+Compared with prior approaches, our framework implicitly captures the long-term temporal structure of behavior through the Bellman equation in the reinforcement learning (RL) formulation.
+The component responsible for encoding long-term and multi-scale dynamics is the motif weight $u(t)$.
+From the linear decomposition of the state–action value function (Eq.
+3),
+$u(t)=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,ds^{\prime},$
+it follows that $u(t)$ integrates two sources of information simultaneously: (i) long-horizon structure through the value function $V(s^{\prime})$, which summarizes discounted future trajectories with an infinite horizon, and (ii) short-horizon structure through the immediate reward parameter $w$.
+Therefore, this formulation naturally enables modeling of multi-scale temporal dependencies.
+Empirically, we observe this structure in Fig.4B: some motifs exhibit stable, slowly varying activation profiles (e.g., motif 1, gray curve), whereas others fluctuate rapidly and capture short-term transitions (e.g., motif 2, pink curve).
+Intuitively, consider a mouse that intermittently sniffs throughout the day but engages in fast running only during a brief morning period.
+The coefficient $u_{i}(t)$ for sniffing would rise gradually across long time windows, while the coefficient $u_{j}(t)$ for running would increase only during short, specific intervals.
+Appendix L Hyperparameter Discussion
+L.1 Motif dimension
+To evaluate the sensitivity of our model to the dimension of the hidden state, we change the dimensions of motifs $d$ and verify them in different datasets.
+In the discrete case, there exists a minimum dimension that could fully represent the motif space.
+In fact, the motif discovery (Eq.
+2) is equivalent to SVD of the transition matrix $P(s^{\prime}|s,a)$(Ren et al., 2022).
+Since the rank of the transition matrix is $|\mathcal{S}|$, the ranks of $\phi$ and $\mu$ matrices are both $|\mathcal{S}|$, which is exactly the minimum dimension of $\phi(s,a)$ that could fully represent the motif space.
+This could be clearly seen in the gridworld dataset (Fig.
+8) and labyrinth navigation dataset (Fig.
+9).
+For the gridworld dataset, $d_{min}=3*3=9$.
+If $d<9$, the SVD could only provide a low-rank approximation.
+As a result, the recovered reward map is blurred and incorrect (Fig.
+8A-C) and the correlation remains low (Fig.
+8G).
+For $d\geq 9$, the motif matrix is full-rank.
+So the reward maps are nearly perfect (Fig.
+8D-F), and the correlations are near 100% (Fig.
+8G).
+Similar analysis also applies to labyrinth navigation dataset where $d_{min}=127$(Fig.
+9).
+In this case, variant with fewer motif dimensions ($d=64$) cannot generate realistic reward maps while variants with more motif dimensions ($d\geq 127$) can faithfully recover the reward peaks in each task.
+Figure 8: Reward maps generated by models of different motif dimension $d$ in the gridworld dataset.
+Figure 9: Reward maps generated by models of different motif dimension $d$ in the labyrinth navigation dataset.
+In the continuous case, since the transition kernel is infinite-dimension, any finite $d$ would only provide a low-rank approximation to the transition kernel.
+As $d$ increases, the fitting performance would be better.
+However, since the importance of different motifs is different, by observing the performance under different $d$, we could select one to be as small as possible while still capturing the entire motif space as comprehensively as possible, and maintain the performance.
+In this experiment (Fig.
+10), we select $d=64$.
+Figure 10: AUC generated by models of different motif dimension $d$ in the animal free-moving dataset.
+L.2 Noise distribution
+The continuous version of MCD is based on contrastive learning.
+Therefore, a high-quality negative sample distribution is crucial for the motif and policy learning.
+Here we evaluate different selections of noise distributions in the continuous free-moving animal behavior dataset.
+For the motif learning (Eq. 7), in the main text, we directly sample states from the dataset. In other words, the noise distribution is $\rho(s)=\int\tau^{e}(s,a)da$.
+In ”Motif-N” variant, we replace it by a uniform distribution, whose dimension-wise bounds are determined as the maximum and minimum values of the states in the dataset, i.e.
+$s^{\prime}_{i}\sim U(\min(\rho(s_{i})),\max(\rho(s_{i})))$.
+Because the negative samples are not good enough, the resulting energy-based model can neither capture high-quality motifs nor provide good basis vectors for later policy learning.
+Therefore the AUC score remains low (Fig 11).
+For the policy learning (Eq.
+8), in the main text, we also sample negative samples directly from the dataset $(s^{\prime},a^{\prime})\sim\tau^{e}$ and use $a^{\prime}$ as a negative sample of actions.
+In other words, the noise distribution is $\zeta(a)=\int\tau^{e}(s,a)ds$.
+In the ”Policy-N” variant, we sample negative pairs from uniform distribution $a^{\prime}_{i}\sim U(\min(\zeta(a_{i})),\max(\zeta(a_{i})))$, similar to ”Motif-N”.
+In the ”AR-Actor” variant, we additionally train an auxilliary autoregressive actor $\pi(a|s)$ to fit $\pi^{e}(a|s)$ using MLE.
+And the negative samples are from this actor $\pi(a|s)$.
+For each dimension, the actor has a network of two hidden layers, with hidden dimension=256 each, to produce the mean and variance of this action dimension, i.e.
+$a_{i}\sim\mathcal{N}(\mu(a_{i}|s,a_{<i}),\sigma^{2}(a_{i}|s,a_{<i}))$.
+As expected, uniform noise distribution is not good enough, and ”Policy-N” achieves lower AUC than the original version.
+And the ”AR-Actor” achieves similar performance on the test dataset, proving the robustness of MCD to noise distribution if it is good enough.
+Figure 11: AUC generated by models trained under different noise distribution in the animal free-moving dataset.
+We also tested the influences of the number of negative samples (Fig.
+12).
+MCD is consistently better than Key-Moseq and maintains a stable performance.
+It turns out that the model is robust against the choice of number of negative samples.
+Figure 12: AUC generated by models trained using different numbers of negative samples for motif contrastive learning in the animal free-moving dataset.
+$nn=k$ means the number of negative samples per positive sample is $k$.
+L.3 Dataset Partition
+In the third experiment, we randomly split the whole dataset into training set and test set.
+Here, we train the model on some mice and then test it on other heldout mice, i.e.
+split the dataset by mouse identities.
+This experiment shows whether the motifs learned by the model are general and transferrable across different subjects.
+The result is shown in Fig.
+13 where the complete exclusion of one animal from the training set does not impair the performance.
+This experiment proves that the MCD can learn a general set of motifs that could be transferred across animals.
+Figure 13: AUC generated by models trained using different train-test split methods.
+For MCD(random), the dataset is randomly split.
+For MCD(heldout), the behavior of one mouse is specially held out as the test set while the rest is taken as the training set.
+L.4 Gaussian random walk prior
+In the free-moving animal behavior dataset (Sec.
+4.3), we impose a Gaussian random-walk prior on the time-varying motif weights $u(t)$ to encourage temporal smoothness.
+To assess the sensitivity of our method to this regularization, we varied the strength of the Gaussian prior and re-evaluated model performance (Fig.
+14).
+As expected, very large coefficients introduce slight degradation due to oversmoothing.
+However, across all tested values, performance remains substantially higher than the baseline model.
+This demonstrates that MCD is robust to the choice of Gaussian random-walk regularization strength and does not rely on fine-tuning this hyperparameter to achieve strong results.
+Figure 14: AUC generated by models tested using different Gaussian random walk prior coefficients in the animal free-moving dataset.
+Appendix M Scientific Impact and Relations to Neuroscience and Ethology
+Scientific relevance to neuroscience.
+Our work is primarily motivated by the neuroscience perspective, where the goal is to extract interpretable behavioral structure that can be directly linked to neural circuit dynamics, internal motivational states, and decision-making processes.
+From this viewpoint, MCD provides scientifically meaningful variables: it discovers low-level motor motifs that correspond to reproducible movement primitives and models behavior as smoothly varying mixtures of these primitives.
+This mirrors known neural control principles in the motor cortex, basal ganglia, and brainstem, where overlapping action components, not discrete switches, combine to generate natural movement.
+The resulting motif representations and time-varying policy weights offer a rich, biologically interpretable representation for analyzing how neural populations evolve alongside behavior.
+The learned representations can be mapped to neural representations in future work.
+Each motif $\phi(s,a)$ defines a low-level motor primitive, giving a clear behavioral regressor for examining whether neurons in motor cortex encode specific movement components.
+The time-varying policy weights $w(t)$ and reward-related weights $u(t)$ describe how motifs are combined and modulated over time, offering hypotheses about potential control- and value-related signals in the dorsal striatum.
+While we do not perform neural analyses in this paper, the structured motif representation provides a principled framework for relating behavior to activity in these circuits in future neuroscience studies.
+For the training protocol, our method is trained purely from behavior; no neural signals, reward labels, or joint behavior–neural objectives are used.
+This distinguishes our approach from models that rely on neural data to infer latent goals or value functions.
+On the other hand, although our method does not require neural recordings, the learned motifs correspond to low-dimensional dynamical components of behavior that can in principle be aligned with neural manifolds (e.g., cyclic modes for gait, motor primitives, or population attractors).
+Relevance to ethology and real-world behavior.
+Beyond neuroscience, MCD is well suited for ethological studies where the goal is to characterize behavior in naturalistic, minimally constrained environments across multiple timescales.
+By capturing long-range dependencies, allowing motifs to co-occur, and modeling continuous dynamics, MCD can describe multi-scale organization of behavior, such as exploratory sequences, foraging patterns, or grooming hierarchies, without assuming discrete states.
+This enables ethologists to quantify how natural behaviors are composed, how they transition, and how they evolve over long durations.
+Thus, while our emphasis is on neuroscience applications, the method is fully compatible with ethological framework and supports general scientific questions about the structure and function of natural behavior.
+Appendix N Generated Rollout Trajectories
+In the animal free-moving dataset, we train a generative energy model to estimate the state–action value function.
+Here, we assess its reliability using Hamiltonian Monte Carlo (HMC) sampling.
+For each motif $d_{i}$, we construct a one-hot vector $u$ with $u_{d_{i}}=1$ and $u_{d_{j}}=0$ for all $j\neq i$, and then sample an action $a$ from the learned energy model.
+For each $d_{i}$, we sample $10$ actions in total.
+The resulting motif-specific rollout trajectories (Fig 15) largely align with the empirical averages (Fig.
+7B), demonstrating that the model captures the dominant behavioral tendencies.
+A small number of trajectories deviate from the empirical trends, probably because $u$ is never used as a strict one-hot vector during training but instead appears as a mixture over motifs.
+We attribute these deviations to unavoidable generative noise and the distributional mismatch between test-time one-hot inputs and the mixed representations observed in real data.
+For that, we choose several clips and generate the rollouts (number of rollout steps=5) based on the inferred $u(t)$.
+The alignment between generated trajectories and real trajectories (Fig.
+16) show that MCD could capture the real behavior dynamics.
+HMC parameters.
+We use a step size of $1\times 10^{-4}$, 100 leapfrog steps per action, temperature =1=1, mass matrix $M=I$, and initial momentum $r\sim\mathcal{N}(0,1/200)$.
+The leapfrog integrator is employed to improve sampling accuracy.
+Figure 15: Motif-specific rollout trajectories generated by HMC sampling from the trained energy model. Indexes are consistent with Fig. 7.
+Figure 16: Rollout trajectories based on inferred $u(t)$ generated by HMC sampling from the trained energy model.